### v1

# Explaining GPT-4’s Schema of Depression Using Machine Behavior Analysis

Abstract
[…]
keywords:
Depression, Language Models, Machine Behavior, GPT-4
1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT-4 or Google’s Gemini are starting to be actively used by thousands of individuals for mental health support, although we have a limited understanding of their outcomes and limitations. As of January 2024, the ChatGPT store [1] reported tens of thousands of conversations involving GPT therapy bots [2], and approximately 1 in 25 chatbot users seek mental health support as one of their primary uses of LLMs [3]. A recent survey in the Journal of Medical Economics revealed that over 1 in 10 clinicians in the USA use ChatGPT, and roughly 50% are interested in continuing its use for data entry, medical scheduling, or research [4].
The potential benefits of LLMs in the mental health domain are numerous. Mental health care often begins with psychodiagnostic assessment, where employing LLMs could play a crucial role for more timely and effective screening and assessment [5]. Present-day screening tools have limited sensitivity, standardization [6], and cultural adaptability [7, 8], plus people report increased willingness to use online mental health technologies available to them [5]. Further, the ability to administer assessments earlier in the care cycle can address the fact that diagnosis often lags behind the onset of symptoms by over a decade [9, 10, 11]. Earlier detection and treatment could mitigate financial barriers that arise from treating more severe conditions [12, 13, 14]. However, similar to treating mental disorders with unevaluated drugs, unevaluated use of LLMs for mental health support comes with many risks [15, 16].
Figure 1: Overview of Machine Behavior Evaluation to uncover the structure of depression in GPT-4. (1) GPT-4 is prompted with step-by-step instructions on performing depression assessment on human-written essays. (2) GPT-4 carried out assessments on data from Gu et al., 2023 [17], containing depression essays written by individuals. The model estimated the severity scores for PHQ-9 symptoms explicitly mentioned in the essay, followed by scoring the remaining symptoms that were implicit in the essay. (3) GPT-4’s estimated scores and spans are used to infer its internal structure of depressive symptoms, which revealed how the web of symptoms was connected, and what language GPT-4 reliably used to estimate depression severity.
This paper presents a framework for evaluating LLMs’ “representation” of mental health. To do this, we apply principles from measurement theory [18] to infer GPT-4’s schema of depression (how it appears to organize, interpret, and relate depressive symptoms). In the process, we quantitatively evaluate two fundamental questions: (a) how GPT-4 identifies and estimates depression symptoms from language, and (b) its latent structure between the symptoms used to form an overall assessment. We utilize this framework to study GPT-4’s structural representation of depression for several reasons. First, depression is one of the most prevalent mental health conditions, affecting an estimated 10–15% of the population throughout their lifetime [19, 20], and is the leading cause of disability worldwide [21]. Second, there exists a vast body of research on depression’s symptoms, structure, and their inter-relationships, providing a strong theoretical reference against which to compare [22, 23, 24]. Lastly, depression is a highly heterogeneous disorder comprised of cognitive, affective, behavioral, and somatic symptoms [25]; depression is also highly co-morbid with other conditions, including anxiety [26], substance use [27], somatic disorders [28] and other internalizing conditions [29]. Thus, understanding GPT-4’s schema of depression suggests generalizations to many aspects and forms of psychopathology.
While previous work on how GPT-4’s assessments are formed is limited, recent studies have evaluated GPT-4’s ability to predict mental disorder [30]. Empirical approaches that seek to explain GPT-4’s behavior are often limited to performance benchmarks and comparative studies [31], which measure the agreement between model’s responses and those of humans [32, 33] or experts [34]. Such methods can reveal what the model can do but do not explain how it performs such tasks. The responsible use of LLMs, particularly in critical applications that involve human health and well-being, requires understanding how they perform tasks —– questions that have traditionally fallen within the purview of fields like psychology and cognitive science.
Over the years, psychology has developed and continually refined tools and methods to measure, interpret, and understand the human mind using mathematical and empirical methods designed with the intent of explaining human mental functioning through behavioral observation [35]. Such methods have uncovered schemas — the internal concepts underlying a process — embedded within the neural system from observations of human behaviors [36, 37, 38]. Drawing parallels from cognitive and behavioral psychology, our experiments are designed to collect extensive observations of GPT-4’s ability to assess depression. We analyze its generated text (responses) and the input prompts (stimuli) to infer GPT-4’s schema of depression. This inferred schema is used to explain GPT-4’s behavior in assessing depression [39]. To the best of our knowledge, this is the first work to perform machine psychology [40] to explain GPT-4’s functioning. We employ a robust methodological framework that can be extended to other LLMs to understand how psychopathology is represented within them. This is depicted in Figure 1 and described extensively under Methods (§4).
The unique contributions of this study center on explaining GPT-4’s behavior when identifying depression risk within open-ended essays. In a sample of 955 individuals, we employ measurement theory to infer GPT-4’s schema of depression by analyzing (a) the relationships among key depression symptoms and (b) the role of each symptom in determining the degree of depression within the language model. This inferred schema is compared with one derived from self-reported scores to quantitatively identify similarities and differences in the conceptualization of depression as a network of interconnected symptoms. We used multivariate linear regression to assess how GPT-4 integrates explicitly mentioned factors in essays to estimate risks for unmentioned factors. We then applied differential language analysis [41, DLA] to identify the linguistic cues associated with GPT-4’s identification of depressive risk factors. Lastly, we evaluated whether GPT-4 shows evidence that it uses the phrases that it cited as explanations for producing its depression scores.
2 Results
GPT-4 was used to analyze language samples from 955 participants, who described their feelings of depression, with the goal of estimating symptom-level scores based on the PHQ-9 [42] questionnaire. This process involved two key steps. In the first step, the model identified symptoms explicitly mentioned in the participants’ essays, citing the corresponding text spans and assigning severity scores to each explicitly mentioned symptom. In the second step, the model inferred any remaining PHQ-9 symptoms that were not directly mentioned in the essays but could be implied from the context, providing a free-form rationale for each inferred symptom, along with a severity estimate.
To validate these estimates, participants also completed the PHQ-9 questionnaire, which assessed their symptom severity over the previous two weeks. Additionally, a random subsample of 209 participant responses was manually annotated by two psychology experts using a similar annotation process to that given to GPT-4. The comparison between GPT-4’s estimates and the participants’ self-reported PHQ-9 scores was conducted on the full dataset, while the comparison between GPT-4 and expert annotations, as shown in Table 1, was based on the subset of expert-annotated data.
[…]
Experts - GPT Self-report - GPT Experts - Self-report
Symptom $r$ $r$ $r$
PHQ 9 Total .81 .70 .68
Anhedonia .65 .59 .56
Depressed .81 .69 .66
Sleep .69 .54 .46
Fatigue .68 .59 .53
Appetite .49 .43 .43
Worthlessness/Guilt .70 .50 .53
Concentration .62 .48 .45
Psychomotor .31 .23 .14
Suicidal Thoughts .63 .31 .30
Figure 2: Similarities and differences between GPT-4’s and self-report’s symptom structure as symptom covariances.
(a) Pearson correlation between symptoms estimated by GPT-4 (left) and self-reported by participants (right). The average correlations of each symptom with rest are presented as column vectors for both GPT-4 and self-report.
(b) Difference between GPT-4’s and self-reports’ Symptom-Symptom correlations. 95% confidence interval of the difference in correlations from bootstrapped resampling over 500 trials is enclosed within brackets.
Convergent Validity: GPT-4 estimates correlate with experts-judgments and self-report.
Table 1 shows the average Pearson correlation ($r$) between GPT-4 estimates and experts for each symptom, as well as the correlation between GPT-4 and self-reported scores, computed on the subset assessed by the experts ($N=209$). GPT-4 demonstrated high agreement with self-report ($r=0.70$) and experts-judgments (avg $r=0.81$) on total PHQ-9 ratings. Across all symptoms, the assessors of depression essays (i.e., GPT-4 and the experts) demonstrated higher agreement with each other than either did with self-reported PHQ-9 scores. The absolute difference between GPT-4’s agreements with self-report and experts (difference between first two columns) was less than $0.10$ for all somatic symptoms (fatigue, appetite, and psychomotor agitation/retardation), except for issues related to Sleep ($\Delta=0.15$). All cognitive and emotional symptoms except for anhedonia had a difference in correlation scores over 10 Pearson points. Highest differences were seen with suicidal ideation ($\Delta=0.32$), worthlessness/guilt ($\Delta=0.20$) and concentration difficulties ($\Delta=0.14$). GPT-4’s agreement with experts ($r_{exp1}=0.80,r_{exp2}=0.81$) on total PHQ-9 score was very similar to the agreement between both the experts ($r=0.83$).
Symptom-Symptom Relationships: GPT-4 and Self-report have similar inter-symptom covariates overall, except for two symptoms.
The symptom-symptom associations from GPT-4 estimates as depicted by the heatmap (left) in 2(a) had positive Pearson correlations across the board ($r=0.23-0.78$). This was in agreement with the symptom-symptom associations from self-reported (right) scores ($r=0.38-0.78$). The effect sizes of the symptom-symptom relationships from GPT-4 estimates were similar to self-report for all but two symptoms, namely, psychomotor agitation/retardation and suicidal ideation. 2(b) clearly shows this difference between symptom-symptom associations from GPT-4 and self-report is statistically significant (95% CI doesn’t include 0.0). GPT-4-estimated severities of psychomotor agitation/retardation had a stronger association with other symptoms than self-report. In contrast, suicidal ideation was less related to other symptoms for GPT-4 than in the self-report. This is evident from the significantly lower association of suicidal ideation with the rest of the symptoms than what was observed in self-reported scores.
Table 2: Symptom’s item location ranking according to IRT over GPT-4 and self-report scores for zero and non-zero decision boundary. A lower rank value corresponds to the symptom being best at distinguishing lower values of depression, while a higher rank indicates being better at distinguishing higher values of depression. Ranks are accompanied by a 95% confidence interval estimated using bootstrapped resampling across individuals (N=955, num. trials=500).
PHQ9 Item Item Location Rank
GPT-4 Self-Report Item location difference
Anhedonia 3 [1, 3] 3 [3, 3] 0
Depressed Mood 1 [1, 2] 2 [1, 2] -1
Sleep 4 [4, 5] 4 [4, 5] 0
Fatigue 2 [1, 3] 1 [1, 2] 1
Appetite 8 [7, 8] 7 [7, 7] 1
Worthlessness / Guilt 6 [6, 6] 5 [4, 6] 1
Difficulty concentrating 5[4, 5] 6[4, 6] -1
Psychomotor 7[7, 8] 8[8, 8] -1
Suicidal Ideation 9[9, 9] 9[9, 9] 0
Differential Item Functioning Analysis: GPT-4 and self-report have similar discriminant characteristics over latent depression scores.
We utilize Item Response Theory (IRT) to conduct a Differential Item Functioning (DIF) analysis on two sets of scores — self-report and GPT-4 estimates. The DIF ranks highlight the depression severity ranges where each item most reliably identifies the presence of its symptom. [43, 44, 45]. A lower-ranked items are best at distinguishing lower latent scores, and higher-ranked items for higher latent scores. A positive difference in an item’s rank between GPT-4 and self-report means that language-based assessment of this item by GPT-4 is more informative than self-report scores to capture higher depression scores. On the other hand, a negative difference implies that the language-based score of this item has lesser discriminant power than the self-report over higher depression scores. Table 2 shows the DIF ranks for GPT-4 and self-report along with its 95% confidence interval estimated through bootstrapped resampling across individuals over 500 trials.
Overall, the distinguishing characteristics of items were in agreement between the two, with suicidal ideation, appetite, and psychomotor agitation/retardation having higher ranks (7-9), and anhedonia, depressed mood, and fatigue having lower ranks (1-3). All items had a very small 95% confidence interval, showing that these rank estimates have high reliability. Suicidal ideation (rank=9), sleep (rank=4), and anhedonia (rank=3) were the three items with no difference in ranks between GPT-4 and self-report. The difference in ranks of other items doesn’t indicate that language-based severity scores of GPT-4 possessed a significantly different characteristic compared to self-report, i.e. there is a high similarity in the item functioning of self-report and GPT-4 for PHQ-9.
Table 3: Contrast of explicit and implicit symptom mentions and convergence. Pearson Correlation ($r$) between GPT-4 and self-reported scores for all symptoms under conditions where symptoms were identified to be explicit and implicit (N=955) by GPT-4. The bar chart in % Explicit column represents the rate of explicit mentions according to GPT-4. Higher correlations were observed consistently when symptoms were identified to be explicitly mentioned.
Symptom Explicit $r$ Explicit vs. % Implicit Implicit $r$
Anhedonia .71 53.72% .43
Depressed Mood .70 96.44% .44
Sleep .52 11.31% .43
Fatigue .58 25.13% .51
Appetite .47 5.03% .33
Worthlessness / Guilt .57 24.19% .45
Concentration .50 9.42% .45
Psychomotor .33 3.35% .24
Suicidal Ideation .75 5.55% .21
Mean .57 26.02% .39
Explicit & Implicit Symptoms: GPT-4 is more precise in estimating explicit symptoms than implicit ones.
Convergence of GPT-4 with self-report in samples identified with explicit mentions of symptoms was higher than unmentioned samples across all symptoms with an average $\Delta$ of +0.18 Pearson points (Table 3). Particularly, GPT-4’s estimation was significantly more ($\Delta\geq 0.25$ Pearson points) in agreement with self-report under explicit mention than implicit for anhedonia, depressed mood, and suicidal ideation. Language describing these three symptoms led to more precise GPT-4 estimates than language describing other symptoms, despite their varying rates of explicit mentions ($range=5.55\%-96.44\%$). The cardinal symptoms, ie., depressed mood and anhedonia, had notably higher rates of explicit mentions in the essays ($96\%$ and $53\%$ respectively) than other symptoms.
Figure 3: Effect of explicit symptom scores on implicit symptom scores as estimated by GPT-4. Wider bands represent larger effect sizes ($\beta$ values, i.e., coefficients of multivariate linear regression from predicting estimated implicit symptom score using all explicit symptom scores while holding other implicit symptoms constant). The 90% confidence interval of the regression coefficients was estimated using bootstrapped resampling over 500 trials. Confidence intervals that included 0.0 were dropped. Explicit refers to the symptoms estimated by GPT-4 when it identified explicit mention in essays and Implicit refers to the symptom scores when they were unmentioned in essays. Dotted bands represent a negative association.
Figure 4: GPT-4’s Language of Depression: Words from the human-written essays that were associated with being identified as explicit symptomatic markers (left) and higher severity score (right) as estimated by GPT-4. Positive Cohen’s d values are concordant with GPT-4-identified explicit markers $(p<0.05)$. Higher Pearson $r$ maps to higher depression severity and words with significant correlations $(p<0.05)$ are boldfaced. Both tests were Bonferroni corrected.
Interplay between explicitly mentioned symptoms and inferred symptoms
We examined how explicit symptoms influence implicit symptoms in GPT-4’s assessments by calculating the regression coefficients ($\beta$) from a multivariate linear regression model. This linear model predicted implicit symptom scores based on explicit symptom scores while controlling for all other implicit symptoms. Figure 3 depicts the effect sizes of explicit symptoms in influencing GPT-4’s estimate of implicit symptoms.
[…]
Except for fatigue, all other symptoms that were explicitly mentioned more than 10% of the time —– including the cardinal symptoms and feelings of worthlessness/guilt —– had a widespread influence on estimating the severity of other unmentioned symptoms. The explicit mention of nearly all the items predicted fatigue. However, fatigue was a very weak indicator of other unmentioned symptoms, despite being explicitly mentioned in 25% of the cases. We identified three pairs of symptoms with strong mutual relationships, where the explicit presence of one symptom influenced the implicit estimation of the other, and vice versa. These pairs include anhedonia and depressed mood, fatigue and psychomotor agitation/retardation, and anhedonia and concentration difficulties.
Uncovering GPT-4’s language of depression: GPT-4 generated reliable reasons for diagnosis
From the left bar chart in Figure 4, it is conclusive that GPT-4 identified explicit markers (positive Cohen’s $d$) of depressive symptoms included words that have been known to be associated with depression from prior studies [46, 47], such as affective words (depressed, feeling low, sad, down), verbs (do), and somatic descriptors (energy, tired, sleep). On the other hand, words from non-spans (negative Cohen’s $d$) mostly consisted of functional words and a few nouns (e.g. pandemic, home). All these words on the left bar chart of Figure 4 had a statistically significant association ($p<0.05$; tested for multiple hypotheses using Bonferroni correction).
In the right bar chart, 12 out of the 25 words with positive Cohen’s $d$ had statistically significant Pearson correlation ($r;p<0.05$) with GPT-4’s total estimated PHQ-9. Essays containing ‘not been depressed’ and mentions of ‘happy’ showed a strong negative association with the total score, and phrases like ‘have been depressed’, ‘sleep’, and ‘low’ (as in low energy) had a positive association. From the set of words with negative Cohen’s $d$, only ‘it’ had a significant but weak association ($r=0.09$) with the total estimated PHQ-9. This showed that GPT-4 doesn’t strongly depend on any meaningful linguistic signal outside its cited explicit markers.
3 Discussion
Receiving professional mental health support is a multi-step, non-linear process that begins with screening and initial assessment. A significant percentage of individuals do not receive timely support due to social barriers and infrastructural limitations within the traditional mental healthcare system [48, 49]. Though effective mental health treatments exist, most individuals with mental disorders do not receive treatment, in part due to shortages of mental healthcare providers [50, 51]. Digital technologies like GPT-4 which possess unprecedented human-comparable skills and may soon be able to practically address many such challenges. With current consumer costs being around \$30/month [52], models such as GPT-4 may be well-positioned to mitigate well-established inhibitors such as high costs, limited mobility, inadequate insurance coverage, and lack of support for parents [15]. Additionally, it can overcome infrastructural limitations, including the need for cultural adaptation and the dearth of mental health clinicians. However, overcoming these challenges does not guarantee precise assessments due to limitations in sensitivity with current measurement tools. To enhance accuracy, these technologies can leverage passive and quasi-passive indicators from daily life [53, 54] to form more precise assessments and aid the delivery of timely care to those in need.
While studies have focused on LLMs’ generations to identify and measure desirable behaviors for mental health support [55, 56], the current study instead sought to inform how the internal structure of depression is represented within an LLM by analyzing its behavior with respect to symptoms. Hence, we tasked GPT-4 with identifying the key symptomatic markers of depression expressed within human essays. These essays contained exhibits of causes (e.g. substance use, sedentary lifestyle) and effects (e.g. tired, exhausted) of depression through direct or synonymous mentions (e.g. “I don’t like doing things anymore” vs “I don’t see the point of getting out of the bed anymore” both suggest lack of interest). From GPT-4’s assessments on language describing depression, we learn how it views the condition to manifest in people, akin to how practicing therapists understand conditions based on the way patients describe it [57, 58].
We evaluated the viability of GPT-4 for assessing depression using two key principles from measurement theory: reliability and validity. We found that GPT-4 displays moderate-high levels of concurrent validity and internal consistency (reliability) with depressive severity from a well-established rating scale measure, PHQ-9. The overall result of GPT-4’s convergence with self-report was consistent with overall associations of GPT-4 [59] and other LLM[33] with standard depression measures, while the item-level results extend convergence to detection of depression symptoms. Further, we found that GPT-4’s generated explanations were generally faithful [60], with little evidence that it relied on linguistic cues outside its explanations of explicit phrases used to estimate PHQ-9 scores. These findings indicate that GPT-4 can be useful at screening patients and identifying symptoms from their language. Language is an information-rich medium compared to rating scales [61] which can provide avenues for making hypotheses about predisposing, precipitating, and perpetuating factors of psychopathology for a given patient while estimating the symptom severity. Thus, the explanatory output of LLM assessment may free up clinicians to prioritize complex case conceptualization and the creation of carefully tailored treatment plans.
[…]
Highly depression-specific symptoms such as anhedonia, depressed mood, and feelings of worthlessness/guilt [22, 62] had both higher base rates of explicit mentions (Table 3) and higher effects on implicit symptoms (Figure 3), explaining the severity estimations when symptoms were unmentioned in the essays. Symptoms with lower levels of specificity to depression [22], like sleep and appetite, were inferrable in the presence of almost any other symptom. These nonspecific symptoms are shared by depression and anxiety-related disorders [22].
Another important consideration in explaining LLMs is the instructions it is given. Given that the information present in the instructions can affect the performance of these models [63, 64, 65, 66], we instructed GPT-4 to estimate symptoms of depression before estimating the severity of overall depression. This provided a means to understand the cohesiveness of the structure of depression, its faithfulness to what it identifies as explicit markers, and varying difficulty levels in estimating the symptoms. This mirrors studies of LLMs in other domains that have demonstrated improved effectiveness when allowed to break a task into compositional steps [67, 68] or when instructed to carry out specific modules that culminate into a task [69, 70].
Using measurement theory to arrive at explanatory schemas can inform how LLMs function, where they are bound to fail, and how to effectively situate them in the mental health support system to maximize overall benefits. There has been increased advocacy for transparency and explainability [71] of LLMs to improve public faith and understand the risks associated with its usage, especially in areas such as mental health support [72]. Our evaluation framework is intended to help clinicians and decision-makers gain insights about model capabilities and pitfalls so that they can determine the capacity at which contemporary models can be used in the current care pipeline and further improve future systems to yield stronger outcomes.
Our work had several limitations. Foremost, the scope of our results is limited to explaining GPT-4’s behaviors when used as a depression assessment tool. Although other modalities of behavior beyond the patient’s verbal behavior can be used for the assessment, we employ language because it intrinsically contains more information, and the advent of LLMs enables us to leverage the power of language to deliver accurate assessments [61, 73] Further exploration is required to assess whether these models are suitable for use in other stages of the mental health treatment pipeline. It is a multi-stage process, encompassing assessment of several other conditions from time to time, and folding in interventions targeted at addressing the causes of patient’s issues. While the nature of the collected data only allowed us to perform a cross-sectional analysis on a small population predominantly from the United Kingdom, we should also evaluate how these models assess mental health for other random samples of the population. Accurate mental health assessments would require taking longitudinal symptom trajectories into account as well [35] to overcome the limitations of cross-sectional evaluations, which may fail to capture the dynamic nature of symptom progression and individual variability over time. Although transformer-based models have been effective in encoding longitudinal representations of behavioral  [74, 75] and heath markers [76], it is vital to understand how LLMs can effectively disentangle both these trajectories over time to assess the mental health severity.
Altogether, our analysis of GPT-4 in assessing depressive symptoms highlights two key insights: (a) moderate to high convergence with self-reported data and (b) an underlying symptom schema that explained its assessments. GPT-4 ’s schema closely matched self-report for major symptoms, but notably downplayed the link with suicidal ideation and overemphasized psychomotor agitation/retardation. This divergence was somewhat mitigated when suicidal ideation was explicitly mentioned, yet the strong overemphasis on psychomotor symptoms led to weaker alignment with both self-reports and expert ratings. Our findings indicate that GPT-4 has the potential to assist clinicians by efficiently assessing patients and prioritizing cases based on depression symptom severity, optimizing time for those in urgent need. Expanding these methods to other mental health conditions and large language models will deepen our understanding of the benefits and limitations of these tools in enhancing mental health care.
4 Methods
Data
Collection. Our data was collected as part of another study [17]. From 1040 participants in total, 258 (24.8%) screened participants reported having been diagnosed with Major Depressive Disorder (MDD) and 259 (24.8%) with Generalized Anxiety Disorder (GAD); and 491 (47.2%) participants were recruited from an unscreened population. 32 (3.2%) respondents were excluded for not answering the control items (see Measures) correctly. Removing participants that failed attention check and incomplete survey responses, the dataset comprised 963 participants, 571 identified as female, 388 as male, 2 preferred not to say, and 2 did not respond. The average age was 33.3 (SD = 11.2; range = 18 – 77) years. Participants reported being from the U.S. (n = 284) and the U.K. (n = 679). Table S1 shows the summary statistics of the self-reported score for each PHQ-9 symptom in our data.
The Swedish National Ethics Review Board deemed the research study (Ethics Application 2020-00730) exempt from requiring ethical approval according to Swedish Law (see §§ 3-4 of the Act [2003:460] on ethical review of research involving humans in Sweden). The research was carried out according to the Declaration of Helsinki. Participants were recruited from the online platform Prolific [77]. The sample comprised a mixture of screened and unscreened respondents.
Measures. Below we only describe the measures used in this study – for all details regarding the data please refer to Gu et al., 2023 [17]. Open-ended question on depression - questions asking participants to describe their depression was adapted from previous research[78] and validated in Gu et al., 2023 [17] (under review). The questions read:
Over the last 2 weeks, have you been depressed or not? Please answer the question by typing at least a paragraph below that indicates whether you have been depressed or not. Try to weigh the strength and the number of aspects that describe if you have been depressed or not so that they reflect your overall personal state of depression. For example, if you have been depressed, then write more about aspects describing this, and if you have not been depressed, then write more about aspects describing that. Write about those aspects that are most important and meaningful to you. Write at least one paragraph in the box.
Following this, participants were administered PHQ-9 [42], a 9 item questionnaire used to assess symptoms of major depressive disorder as defined in the DSM-IV [57]. The rating scale included nine items targeting symptoms such as “Feeling down, depressed, or hopeless”, referring to experiences from the last two weeks. A four-point rating scale is used (0 = Not at all to 3 = Nearly every day) to indicate the experienced severity of each symptom.
Procedure. Participants were first given information about the study, including information that participation was voluntary, that they could withdraw at any time, and that their answers were anonymous. Before starting the study, all participants consented to take part in the study and to share their anonymized data openly. The survey had three parts: First, participants answered open-ended questions concerning depression and anxiety/worry, which were presented in random order across participants. Second, they filled out the rating scales, which were randomized. Last, they reported the demographics, and mental health-related questions like sick leave and healthcare visits. The median time to complete the study was 20.1 minutes (mean=24.0, SD=12.8).
Figure 5: Instruction to GPT-4 with an example input text. The instruction given to GPT-4 for depression assessment comprised three parts: a brief task description, detailed step-by-step instructions, and the formatting structure. GPT-4 was prompted with individual depression essays along with this instruction.
Your task is to conduct an in-depth analysis of a provided written text, with the goal of mirroring the psychological state of the author to accurately fill out the Patient Health Questionnaire (PHQ9). The PHQ9 consists of 9 items that are designed to identify symptoms of depression: Anhedonia: Little interest or pleasure in doing things Depressed Mood: Feeling down, depressed, or hopeless Insomnia or Hypersomnia: Trouble falling or staying asleep, or sleeping too much Fatigue: Feeling tired or having little energy Poor Appetite or Overeating Worthlessness or Guilt: Feeling bad about yourself or that you are a failure or have let yourself or your family down Difficulty Concentrating: Trouble concentrating on things Psychomotor Agitation or Retardation: Moving or speaking so slowly that other people could have noticed, or being so fidgety or restless that you have been moving around a lot more than usual Suicidal Ideation: Thoughts that you would be better off dead, or of hurting yourself The text should be scrutinized carefully and the following steps should be strictly followed: STEP 1: Pinpoint and record only the PHQ9 items that are clearly stated in the text, assigning each a severity score between 0–3. At this stage, refrain from evaluating or recording items that are not explicitly mentioned. Enclose the text spans indicative of respective symptoms inside single quotes. If multiple spans are present, separate them using comma. STEP 2: For the remaining PHQ9 items not explicitly mentioned in the text, assign an estimated score. This estimation should be based on what you would expect it to be while factoring the severity score of the symptoms already identified. STEP 3: Combine the scores from the previous steps and present the total PHQ9 score as “Combined Score:”. Categorize this score into only one of the following: None, Mild, Moderate, Moderately Severe, and Severe, and present it as “Category:”. This will signify the completion of the task. To ensure clarity and easy readability, format your output into a nested JSON. The first level should contain the step number as key (‘STEP #’), and the value should be a JSON containing the items as keys and a list containing reason(s) and the estimated severity score as values. The final step’s JSON should contain ‘Final’ as the key and a list containing the summed score and the evaluated category as the value. Here’s the abstracted format of the JSON, with elements inside the angle brackets being placeholders for the actual values:
    "STEP 1": {
        "<Item a>": ["<Reason>", "<Severity Score a>"],
        "<Item b>": ["<Reason>", "<Severity Score b>"],
    "STEP 2": {
        "<Item x>": ["<Reason>", "<Severity Score x>"],
        "<Item y>": ["<Reason>", "<Severity Score y>"],
    "STEP 3": {
        "Final": ["Combined Score: <score>", "Category: <category>"]
Keep in mind, a score of 0 indicates that the symptom was not experienced, while a score of 3 signifies a high severity level of the specific symptom.
Text: “I feel overwhelmed because I am bearing the weight of all the domestic and financial responsibilities. I should be getting more help than I am and it makes me feel resentful and annoyed with myself for allowing it to continue. I am avoiding talking to people because my emotions are unpredictable and I worry that the mask might slip.”
Figure 6: Example Output of GPT-4. The output from GPT-4 shows the assessment results in a nested JSON format, following the steps outlined in the instruction. Step 1 contains PHQ-9 symptoms identified by GPT-4 to be explicitly mentioned in the depression essay, along with the indicative excerpts and a symptom severity score. Step 2 has all the remaining PHQ-9 symptoms accompanied by a free form reason text and an estimated severity score. Finally, Step 3 has the combined estimated PHQ-9 score and an overall degree of depression category.
STEP 1: { Depressed Mood: [‘I feel overwhelmed’, 2], Worthlessness or Guilt: [‘annoyed with myself for allowing it to continue’, 2], Difficulty Concentrating: [‘my emotions are unpredictable’, 0] } STEP 2: { Anhedonia: [“Based on the overall tone of hopelessness and resentment, it is likely that the author is experiencing diminished interest or pleasure”, 1], Insomnia or Hypersomnia: [“No explicit mention, but stress and emotional unpredictability may affect sleep patterns”, 1], Fatigue: [“Feeling overwhelmed could imply a state of mental fatigue”, 1], Poor appetite or overeating: [“No explicit mention, but stress often affects eating habits”, 1], Psychomotor Agitation or Retardation: [“‘avoiding talking to people’ may suggest a withdrawal that could be associated with psychomotor retardation”, 1], Suicidal Ideation: [“No explicit mention of suicidal thoughts or self-harm”, 0] } STEP 3: { Final: [Combined Score: 9, Category: Moderate] }
Experiment Design
Prompt. Only the text response prompted to describe one’s depression was sent to GPT-4 111versioned 11/06 and referred to as gpt4-1106 in openAI’s GPT directory along with the instruction described in Table 5. The instruction was composed into three parts:
  1. (A)
    Task description: A short paragraph describing the task of estimating the PHQ-9 item scores by analyzing the human-written text. The description included the commonly abbreviated form of the 9 items from the PHQ-9 questionnaire.
  2. (B)
    Task Steps: A detailed step-by-step instructions dictating GPT-4 to first estimate the scores for the items that are explicitly stated in the text, followed by scoring the items that were not explicitly stated and finally, concluding with the combined PHQ9 score and severity category.
  3. (C)
    Output Format: GPT-4 was instructed to format its generation into a predefined data structure (JSON object) to facilitate easy processing of data for analysis.
Although, GPT-4 generated outputs fulfilled the completion of PHQ-9 estimation for all the data, it was not completely faithful to the details of the instructions. There were a few issues observed with following the steps detailed in the instructions, especially step 1. 59 (6.1%) out of 956 samples, had small issues that were manually rectified. These included cases where GPT-4 had
  1. (i)
    made subtle changes to the symptom abbreviations (E.g. ’Feelings of Worthlessness or Guilt’ instead of ’Worthlessness or Guilt’),
  2. (ii)
    identified specific aspects of the symptoms (E.g. ’Insomnia’ in ’Insomnia or Hypersomnia’, ’Psychomotor Retardation’ in ’Psychomotor Agitation/Retardation’, ’Poor Appetite’ in ’Poor Appetite or overeating’ etc.),
  3. (iii)
    split the individual steps into separate JSONs, instead of nesting them within a single JSON as described in the prompt.
  4. (iv)
    included a summary of the assessment which was a short essay of its JSON output
  5. (v)
    included symptoms that were not a part of the PHQ-9 (E.g. Financial Worries, Social Exclusion)
For the first two cases, the symptoms were manually corrected to match the abbreviations provided in Step 1 of the instructions. Type iii generations were rectified by manually reformatting the JSON, and the summaries in type iv generations were removed. The symptoms in type v generations were left out of the analysis.
GPT-4’s generations varied based on its hyperparameters. For this work, the temperature value was set to 0.0 to select the most likely token at each step and to allow reproducibility of experiments. Other hyperparameters include frequency penalty (0.1), presence penalty (0.0), top p (1.0), and max tokens (450).
Cleaning and Filtering GPT-4’s Responses
GPT-4 generated responses which were in JSON format as shown in Figure 6 were turned into tabular data by extracting all the scores into individual columns along with a binary variable indicating whether a symptom was found to be explicit by GPT-4. Table S1 shows the descriptive statistics of the individual symptom scores estimated by GPT-4. The excerpts from the user-written text identified by GPT-4 as explicit symptom markers were extracted using simple regular expressions and stored as a CSV.
Expert Judgements
Two psychology experts independently assessed the depression severity from user-written text on a random sample of 209 instances from the same dataset. They were instructed to read the essay to identify excerpts indicative of PHQ-9 symptoms and estimate a severity score for them. If there were no indicative excerpts, they were asked to estimate the symptom severity based on the essay and the estimated explicit symptom scores. The symptoms accompanied by excerpts would pertain to explicit markers as identified by experts.
Analysis
Item Response Theory. In IRT, data from dichotomous or ordinal items are modeled using one or more latent variables with nonlinear link functions, most commonly logistic functions [79]. For this analysis, we used the 1-PL dichotomous logistic IRT model, aka Rasch model [80] to fit item parameters and assess the presence of each of the nine symptoms. This model constrains the slope parameter of the logistic function to 1.0, with the intercept or item location estimated from the responses. The item location parameter indicates where each item stands in relation to assessing depression. We analyzed the rankings of item locations to compare the two sets of scores across all 955 participants.
Interplay Between Explicit and Implicit Symptoms. For each symptom $i$ from the set of all symptoms $P$ (as defined by the PHQ-9), we trained a linear regression model to predict the GPT-4-estimated score of symptom $i$ when it was not explicitly identified. The model was trained to minimize the sum of squared errors, using the scores of all other explicitly identified symptoms $s_{j}^{exp}$, where $j\in P$ and $j\neq i$, as predictors. The linear regression coefficients $\beta$ represent the association between the implicit score of symptom $i$ and the explicit scores of the remaining symptoms, specifically when symptom $i$ was not explicitly identified by GPT-4 in the user text:
$s_{i}^{imp}=\sum_{\begin{subarray}{c}j\in P\\ j\neq i\end{subarray}}\beta_{j}s_{j}^{exp}$ (1)
In cases where the score for $s_{j}^{exp}$ was not explicit (i.e., not identified by GPT-4), it was replaced by the mean score for that symptom. This process was repeated 500 times using bootstrapped samples of the data. Coefficients whose 90% confidence intervals included 0 were excluded from Figure 3.
Language Of Depression. For each user-written text, 1, 2, and 3 grams usage was computed as a relative frequency for each user-written text split into two groups, ie., tokens that were present in the GPT-4’s excerpt spans and tokens that were left out of the spans. We removed all n-grams used by less than 10 participants and limited them to n-grams that were only composed of alphabets (filtering n-grams that had numbers and special characters). We computed Cohen’s d [81] using n-grams’ relative frequencies when it was found in explicit spans and the rest of the user text. After applying Bonferroni correction, it resulted in 56 tokens associated with explicit indication of symptoms (positive d) and 30 that were implicit (negative d) that were statistically significant ($p<0.05$).
We computed the n-gram (1, 2, and 3) relative frequencies for each user-written text, and created two separate feature sets containing the relative frequencies of positive and negative Cohen’s d respectively. Both features were filtered to n-grams with a Pointwise Mutual Information [82] of at least 4.0, which reduced it to 45 and 25 features respectively. To uncover if GPT-4 estimates rely on any linguistic signals outside the explicit markers it generated, these sets of positive and negative words’ relative frequencies (computed per user) were correlated with the total PHQ-9 score from GPT-4. This analysis was performed using DLATK[83], a Python library for language analysis.
Declarations
[…]
Funding AVG and RLB were supported in part by NIH grant R01-AA028032, VV was supported in part by a DARPA grant #W911NF-20-1-0306, and HAS was supported in part by both the grants. YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290. JCE was supported by Stanford’s Institute for Human-Centered AI. LF was supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence research on LLMs, and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University. ONEK, KK and VCE were supported by FORTE (STY-2022/0007; 2022-01022).
The conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, NIH, any other government organization, or the U.S./ German Government.
[…]
CoI/Competing Interest ONEK and KK co-founded a start-up that uses computational language assessments to diagnose mental health problems.
[…]
Ethics Approval and Consent to participate This study was deemed exempt from IRB approval. All study participants’ data was anonymized and they consented to share the data for research.
[…]
Data availability The data for this project will be made available through Github. All data including the original study data, GPT-4 responses, cleaned version, as well as the processed features will be open-sourced for replication.
[…]
Materials availability All materials including the manuscript, display Items (tables and figures), code, and data will be available to the public through Github.
[…]
Code availability All programs for this work were coded in Python and will be open-sourced with the data in the same Github repository. This includes programs that accessed GPT-4 to generate responses using OpenAI’s API, cleaning the data, and our analysis.
[…]
Author Contributions AVG, LF, and HAS designed the research; AVG, YKL, and VV built the analytic tools; ONEK and KK were involved in data collection using Prolific; Clinical psychology experts VCE and KK carried out the expert judgements on the depression essays; AVG, VV, YKL, ECS, JCE, KK, ONEK, VCE, HAS, RLB and LF performed the analysis and wrote the manuscript; TD and AVG made all the display items; All authors made the final decision to submit the manuscript.
Supplementary
The distribution of severity scores for self-report and GPT-4’s assessments over 955 samples from Gu et al., 2023 [17] is shown in  Table S1. While, GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms, showing a different distribution at higher values. This shows that PHQ-9 displays an under-estimation bias of severity scores.
Table S1: Distribution of self-report and GPT-4 estimated scores for each item on PHQ-9 (N=955). GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, but its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms.
Item % Scoring 0 % Scoring 1 % Scoring 2 % Scoring 3
Self Report GPT-4 Self Report GPT-4 Self Report GPT-4 Self Report GPT-4
Anhedonia 24.3 28.5 32.5 43.4 24.4 26.5 18.8 1.7
Depressed Mood 23.8 25.0 32.1 24.6 22.0 44.0 22.2 6.5
Sleep 25.1 35.2 23.8 54.4 21.1 10.1 30.0 0.4
Fatigue 15.4 29.3 27.2 50.9 22.7 19.4 34.6 0.4
Appetite 30.6 57.2 25.9 40.0 19.1 2.7 24.4 0.1
Worthlessness / Guilt 29.4 43.3 25.4 37.5 23.0 18.3 22.2 0.9
Difficulty concentrating 28.0 36.0 30.9 58.3 24.1 5.7 16.9 0.0
Psychomotor Agi./ Ret. 55.5 53.3 23.2 45.3 15.0 1.4 6.4 0.0
Suicidal ideation 60.3 87.9 21.4 8.8 10.0 2.3 8.3 1.1

### v2

@@
-Explaining GPT-4’s Schema of Depression Using Machine Behavior Analysis
+Explaining GPTs’ Schema of Depression: A Machine Behavior Analysis
 Abstract
-[…]
-keywords:
-Depression, Language Models, Machine Behavior, GPT-4
+Use of large language models such as ChatGPT (GPT-4/GPT-5) for mental health support has grown rapidly, emerging as a promising route to assess and help people with mood disorders, like depression.
+However, we have a limited understanding of these language models’ schema of mental disorders, that is, how it internally associates and interprets symptoms of such disorders.
+In this work, we leveraged contemporary measurement theory to decode how GPT-4 and GPT-5 interrelates depressive symptoms, providing an explanation of how LLMs apply what they learn and informing clinical applications.
+We found that GPT-4: (a) had strong convergent validity with standard instruments and expert judgments (r=.70.81)(r=.70-.81), and (b) behaviorally linked depression symptoms with each other (symptom inter-correlates r=.23.78r=.23-.78) in accordance with established literature on depression; however, it (c) underemphasized the relationship between suicidality and other symptoms while overemphasizing psychomotor symptoms; and (d) suggested novel hypotheses of symptom mechanisms, for instance, indicating that sleep and fatigue are broadly influenced by other depressive symptoms, while worthlessness/guilt is only tied to depressed mood.
+GPT-5 showed a slightly lower convergence with self-report, a difference our machine-behavior analysis makes interpretable through shifts in symptom–symptom relationships.
+These insights provide an empirical foundation for understanding language model’s mental health assessments and demonstrate a generalizable approach for explainability in other models and disorders.
+Our findings can guide key stakeholders to make informed decisions for effectively situating these technologies in the care system.
+Keywords: Depression, Large Language Models, GPT, Machine Behavior
 1 Introduction
-Large Language Models (LLMs), such as OpenAI’s GPT-4 or Google’s Gemini are starting to be actively used by thousands of individuals for mental health support, although we have a limited understanding of their outcomes and limitations.
-As of January 2024, the ChatGPT store [1] reported tens of thousands of conversations involving GPT therapy bots [2], and approximately 1 in 25 chatbot users seek mental health support as one of their primary uses of LLMs [3].
+Large Language Models (LLMs), such as OpenAI’s GPT-4 or Google’s Gemini are starting to be actively used by thousands of individuals [18] for mental health support, although we have a limited understanding of their outcomes and limitations.
 A recent survey in the Journal of Medical Economics revealed that over 1 in 10 clinicians in the USA use ChatGPT, and roughly 50% are interested in continuing its use for data entry, medical scheduling, or research [42].
-The potential benefits of LLMs in the mental health domain are numerous.
-Mental health care often begins with psychodiagnostic assessment, where employing LLMs could play a crucial role for more timely and effective screening and assessment [5].
-Present-day screening tools have limited sensitivity, standardization [6], and cultural adaptability [7, 8], plus people report increased willingness to use online mental health technologies available to them [5].
-Further, the ability to administer assessments earlier in the care cycle can address the fact that diagnosis often lags behind the onset of symptoms by over a decade [9, 10, 11].
-Earlier detection and treatment could mitigate financial barriers that arise from treating more severe conditions [12, 13, 14].
-However, similar to treating mental disorders with unevaluated drugs, unevaluated use of LLMs for mental health support comes with many risks [15, 16].
-Figure 1: Overview of Machine Behavior Evaluation to uncover the structure of depression in GPT-4.
-(1) GPT-4 is prompted with step-by-step instructions on performing depression assessment on human-written essays.
-(2) GPT-4 carried out assessments on data from Gu et al., 2023 [17], containing depression essays written by individuals.
-The model estimated the severity scores for PHQ-9 symptoms explicitly mentioned in the essay, followed by scoring the remaining symptoms that were implicit in the essay.
-(3) GPT-4’s estimated scores and spans are used to infer its internal structure of depressive symptoms, which revealed how the web of symptoms was connected, and what language GPT-4 reliably used to estimate depression severity.
-This paper presents a framework for evaluating LLMs’ “representation” of mental health.
-To do this, we apply principles from measurement theory [18] to infer GPT-4’s schema of depression (how it appears to organize, interpret, and relate depressive symptoms).
-In the process, we quantitatively evaluate two fundamental questions: (a) how GPT-4 identifies and estimates depression symptoms from language, and (b) its latent structure between the symptoms used to form an overall assessment.
-We utilize this framework to study GPT-4’s structural representation of depression for several reasons.
+However, similar to treating mental disorders with drugs that are not understood, there are many risks in using LLMs for mental health support without explanation of how they behave in the domain [45, 8].
+This paper presents a framework for evaluating and explaining LLMs’ “representation” of mental health.
+To do this, we apply principles from measurement theory [37] to infer GPT-4’s and GPT-5’s schema of depression (how it appears to organize, interpret, and relate depressive symptoms) based on its behavior.
+In the process, we quantitatively evaluate two fundamental questions: (a) how does GPT identify and estimate depression symptoms from language, and (b) what is its latent structure between the symptoms used to form an overall assessment.
+We utilize this framework to study GPTs’ structural representation of depression for several reasons.
 First, depression is one of the most prevalent mental health conditions, affecting an estimated 10–15% of the population throughout their lifetime [30, 32], and is the leading cause of disability worldwide [13].
@@
 Thus, understanding GPT-4’s schema of depression suggests generalizations to many aspects and forms of psychopathology.
-While previous work on how GPT-4’s assessments are formed is limited, recent studies have evaluated GPT-4’s ability to predict mental disorder [30].
-Empirical approaches that seek to explain GPT-4’s behavior are often limited to performance benchmarks and comparative studies [31], which measure the agreement between model’s responses and those of humans [32, 33] or experts [34].
-Such methods can reveal what the model can do but do not explain how it performs such tasks.
-The responsible use of LLMs, particularly in critical applications that involve human health and well-being, requires understanding how they perform tasks —– questions that have traditionally fallen within the purview of fields like psychology and cognitive science.
-Over the years, psychology has developed and continually refined tools and methods to measure, interpret, and understand the human mind using mathematical and empirical methods designed with the intent of explaining human mental functioning through behavioral observation [35].
+While previous work on how GPTs’ assessments are formed is limited, there is growing evidence of these models’ ability to predict mental disorder [41] as compared to self-report questionnaires [14] or experts [31].
+Such methods reveal what the model can do but do not explain how (i.e.
+explainability) it performs such tasks.
+The responsible use of LLMs, particularly in critical applications that involve human health and well-being, requires explaining how they perform their tasks.
+To address the complexities of explaining a massive model, we use a machine behavior approach [35].
+Over the years, behavioral science has developed and continually refined tools and methods to explain (measure, interpret, and understand) the human mind using mathematical and empirical methods through behavioral observation [15].
 Such methods have uncovered schemas — the internal concepts underlying a process — embedded within the neural system from observations of human behaviors [48, 39, 54].
 Drawing parallels from cognitive and behavioral psychology, our experiments are designed to collect extensive observations of GPT-4’s ability to assess depression.
-We analyze its generated text (responses) and the input prompts (stimuli) to infer GPT-4’s schema of depression.
-This inferred schema is used to explain GPT-4’s behavior in assessing depression [39].
-To the best of our knowledge, this is the first work to perform machine psychology [40] to explain GPT-4’s functioning.
+We analyze its generated text (responses) and the input prompts (stimuli) to infer GPT-4’s and GPT-5’s schema of depression.
+This inferred schema is used to explain the models’ behavior in assessing depression.
+Rahwan et al., (2019) [35] and Markowitz & Boyd (2024) [33] call for such an approach to explain GPTs’ functioning, and to the best of our knowledge, our study represents the first.
 We employ a robust methodological framework that can be extended to other LLMs to understand how psychopathology is represented within them.
-This is depicted in Figure 1 and described extensively under Methods (§4).
-The unique contributions of this study center on explaining GPT-4’s behavior when identifying depression risk within open-ended essays.
-In a sample of 955 individuals, we employ measurement theory to infer GPT-4’s schema of depression by analyzing (a) the relationships among key depression symptoms and (b) the role of each symptom in determining the degree of depression within the language model.
+This is depicted in Figure 1 and has been described extensively under Methods.
+The unique contributions of this study center on explaining GPT-4’s and GPT-5’s behavior when identifying depression risk within open-ended essays.
+In a sample of 955 individuals, we employ measurement theory to infer GPTs’ schema of depression by analyzing (a) the relationships among key depression symptoms and (b) the role of each symptom in determining the degree of depression within the language model.
 This inferred schema is compared with one derived from self-reported scores to quantitatively identify similarities and differences in the conceptualization of depression as a network of interconnected symptoms.
-We used multivariate linear regression to assess how GPT-4 integrates explicitly mentioned factors in essays to estimate risks for unmentioned factors.
-We then applied differential language analysis [41, DLA] to identify the linguistic cues associated with GPT-4’s identification of depressive risk factors.
-Lastly, we evaluated whether GPT-4 shows evidence that it uses the phrases that it cited as explanations for producing its depression scores.
-2 Results
-GPT-4 was used to analyze language samples from 955 participants, who described their feelings of depression, with the goal of estimating symptom-level scores based on the PHQ-9 [42] questionnaire.
-This process involved two key steps.
-In the first step, the model identified symptoms explicitly mentioned in the participants’ essays, citing the corresponding text spans and assigning severity scores to each explicitly mentioned symptom.
-In the second step, the model inferred any remaining PHQ-9 symptoms that were not directly mentioned in the essays but could be implied from the context, providing a free-form rationale for each inferred symptom, along with a severity estimate.
-To validate these estimates, participants also completed the PHQ-9 questionnaire, which assessed their symptom severity over the previous two weeks.
-Additionally, a random subsample of 209 participant responses was manually annotated by two psychology experts using a similar annotation process to that given to GPT-4.
-The comparison between GPT-4’s estimates and the participants’ self-reported PHQ-9 scores was conducted on the full dataset, while the comparison between GPT-4 and expert annotations, as shown in Table 1, was based on the subset of expert-annotated data.
-[…]
-Experts - GPT Self-report - GPT Experts - Self-report
-Symptom $r$ $r$ $r$
-PHQ 9 Total .81 .70 .68
-Anhedonia .65 .59 .56
-Depressed .81 .69 .66
-Sleep .69 .54 .46
-Fatigue .68 .59 .53
-Appetite .49 .43 .43
-Worthlessness/Guilt .70 .50 .53
-Concentration .62 .48 .45
-Psychomotor .31 .23 .14
-Suicidal Thoughts .63 .31 .30
-Figure 2: Similarities and differences between GPT-4’s and self-report’s symptom structure as symptom covariances.
-(a) Pearson correlation between symptoms estimated by GPT-4 (left) and self-reported by participants (right).
-The average correlations of each symptom with rest are presented as column vectors for both GPT-4 and self-report.
-(b) Difference between GPT-4’s and self-reports’ Symptom-Symptom correlations.
-95% confidence interval of the difference in correlations from bootstrapped resampling over 500 trials is enclosed within brackets.
-Convergent Validity: GPT-4 estimates correlate with experts-judgments and self-report.
-Table 1 shows the average Pearson correlation ($r$) between GPT-4 estimates and experts for each symptom, as well as the correlation between GPT-4 and self-reported scores, computed on the subset assessed by the experts ($N=209$).
-GPT-4 demonstrated high agreement with self-report ($r=0.70$) and experts-judgments (avg $r=0.81$) on total PHQ-9 ratings.
-Across all symptoms, the assessors of depression essays (i.e., GPT-4 and the experts) demonstrated higher agreement with each other than either did with self-reported PHQ-9 scores.
-The absolute difference between GPT-4’s agreements with self-report and experts (difference between first two columns) was less than $0.10$ for all somatic symptoms (fatigue, appetite, and psychomotor agitation/retardation), except for issues related to Sleep ($\Delta=0.15$).
-All cognitive and emotional symptoms except for anhedonia had a difference in correlation scores over 10 Pearson points.
-Highest differences were seen with suicidal ideation ($\Delta=0.32$), worthlessness/guilt ($\Delta=0.20$) and concentration difficulties ($\Delta=0.14$).
-GPT-4’s agreement with experts ($r_{exp1}=0.80,r_{exp2}=0.81$) on total PHQ-9 score was very similar to the agreement between both the experts ($r=0.83$).
-Symptom-Symptom Relationships: GPT-4 and Self-report have similar inter-symptom covariates overall, except for two symptoms.
-The symptom-symptom associations from GPT-4 estimates as depicted by the heatmap (left) in 2(a) had positive Pearson correlations across the board ($r=0.23-0.78$).
-This was in agreement with the symptom-symptom associations from self-reported (right) scores ($r=0.38-0.78$).
-The effect sizes of the symptom-symptom relationships from GPT-4 estimates were similar to self-report for all but two symptoms, namely, psychomotor agitation/retardation and suicidal ideation.
-2(b) clearly shows this difference between symptom-symptom associations from GPT-4 and self-report is statistically significant (95% CI doesn’t include 0.0).
-GPT-4-estimated severities of psychomotor agitation/retardation had a stronger association with other symptoms than self-report.
-In contrast, suicidal ideation was less related to other symptoms for GPT-4 than in the self-report.
-This is evident from the significantly lower association of suicidal ideation with the rest of the symptoms than what was observed in self-reported scores.
-Table 2: Symptom’s item location ranking according to IRT over GPT-4 and self-report scores for zero and non-zero decision boundary.
-A lower rank value corresponds to the symptom being best at distinguishing lower values of depression, while a higher rank indicates being better at distinguishing higher values of depression.
-Ranks are accompanied by a 95% confidence interval estimated using bootstrapped resampling across individuals (N=955, num.
-trials=500).
-PHQ9 Item Item Location Rank
-GPT-4 Self-Report Item location difference
-Anhedonia 3 [1, 3] 3 [3, 3] 0
-Depressed Mood 1 [1, 2] 2 [1, 2] -1
-Sleep 4 [4, 5] 4 [4, 5] 0
-Fatigue 2 [1, 3] 1 [1, 2] 1
-Appetite 8 [7, 8] 7 [7, 7] 1
-Worthlessness / Guilt 6 [6, 6] 5 [4, 6] 1
-Difficulty concentrating 5[4, 5] 6[4, 6] -1
-Psychomotor 7[7, 8] 8[8, 8] -1
-Suicidal Ideation 9[9, 9] 9[9, 9] 0
-Differential Item Functioning Analysis: GPT-4 and self-report have similar discriminant characteristics over latent depression scores.
-We utilize Item Response Theory (IRT) to conduct a Differential Item Functioning (DIF) analysis on two sets of scores — self-report and GPT-4 estimates.
-The DIF ranks highlight the depression severity ranges where each item most reliably identifies the presence of its symptom.
-[43, 44, 45].
-A lower-ranked items are best at distinguishing lower latent scores, and higher-ranked items for higher latent scores.
-A positive difference in an item’s rank between GPT-4 and self-report means that language-based assessment of this item by GPT-4 is more informative than self-report scores to capture higher depression scores.
-On the other hand, a negative difference implies that the language-based score of this item has lesser discriminant power than the self-report over higher depression scores.
-Table 2 shows the DIF ranks for GPT-4 and self-report along with its 95% confidence interval estimated through bootstrapped resampling across individuals over 500 trials.
-Overall, the distinguishing characteristics of items were in agreement between the two, with suicidal ideation, appetite, and psychomotor agitation/retardation having higher ranks (7-9), and anhedonia, depressed mood, and fatigue having lower ranks (1-3).
-All items had a very small 95% confidence interval, showing that these rank estimates have high reliability.
-Suicidal ideation (rank=9), sleep (rank=4), and anhedonia (rank=3) were the three items with no difference in ranks between GPT-4 and self-report.
-The difference in ranks of other items doesn’t indicate that language-based severity scores of GPT-4 possessed a significantly different characteristic compared to self-report, i.e.
-there is a high similarity in the item functioning of self-report and GPT-4 for PHQ-9.
-Table 3: Contrast of explicit and implicit symptom mentions and convergence.
-Pearson Correlation ($r$) between GPT-4 and self-reported scores for all symptoms under conditions where symptoms were identified to be explicit and implicit (N=955) by GPT-4.
-The bar chart in % Explicit column represents the rate of explicit mentions according to GPT-4.
-Higher correlations were observed consistently when symptoms were identified to be explicitly mentioned.
-Symptom Explicit $r$ Explicit vs. % Implicit Implicit $r$
-Anhedonia .71 53.72% .43
-Depressed Mood .70 96.44% .44
-Sleep .52 11.31% .43
-Fatigue .58 25.13% .51
-Appetite .47 5.03% .33
-Worthlessness / Guilt .57 24.19% .45
-Concentration .50 9.42% .45
-Psychomotor .33 3.35% .24
-Suicidal Ideation .75 5.55% .21
-Mean .57 26.02% .39
-Explicit & Implicit Symptoms: GPT-4 is more precise in estimating explicit symptoms than implicit ones.
-Convergence of GPT-4 with self-report in samples identified with explicit mentions of symptoms was higher than unmentioned samples across all symptoms with an average $\Delta$ of +0.18 Pearson points (Table 3).
-Particularly, GPT-4’s estimation was significantly more ($\Delta\geq 0.25$ Pearson points) in agreement with self-report under explicit mention than implicit for anhedonia, depressed mood, and suicidal ideation.
-Language describing these three symptoms led to more precise GPT-4 estimates than language describing other symptoms, despite their varying rates of explicit mentions ($range=5.55\%-96.44\%$).
-The cardinal symptoms, ie., depressed mood and anhedonia, had notably higher rates of explicit mentions in the essays ($96\%$ and $53\%$ respectively) than other symptoms.
-Figure 3: Effect of explicit symptom scores on implicit symptom scores as estimated by GPT-4.
-Wider bands represent larger effect sizes ($\beta$ values, i.e., coefficients of multivariate linear regression from predicting estimated implicit symptom score using all explicit symptom scores while holding other implicit symptoms constant).
-The 90% confidence interval of the regression coefficients was estimated using bootstrapped resampling over 500 trials.
-Confidence intervals that included 0.0 were dropped.
-Explicit refers to the symptoms estimated by GPT-4 when it identified explicit mention in essays and Implicit refers to the symptom scores when they were unmentioned in essays.
-Dotted bands represent a negative association.
-Figure 4: GPT-4’s Language of Depression: Words from the human-written essays that were associated with being identified as explicit symptomatic markers (left) and higher severity score (right) as estimated by GPT-4.
-Positive Cohen’s d values are concordant with GPT-4-identified explicit markers $(p<0.05)$.
-Higher Pearson $r$ maps to higher depression severity and words with significant correlations $(p<0.05)$ are boldfaced.
-Both tests were Bonferroni corrected.
-Interplay between explicitly mentioned symptoms and inferred symptoms
-We examined how explicit symptoms influence implicit symptoms in GPT-4’s assessments by calculating the regression coefficients ($\beta$) from a multivariate linear regression model.
-This linear model predicted implicit symptom scores based on explicit symptom scores while controlling for all other implicit symptoms.
-Figure 3 depicts the effect sizes of explicit symptoms in influencing GPT-4’s estimate of implicit symptoms.
-[…]
-Except for fatigue, all other symptoms that were explicitly mentioned more than 10% of the time —– including the cardinal symptoms and feelings of worthlessness/guilt —– had a widespread influence on estimating the severity of other unmentioned symptoms.
-The explicit mention of nearly all the items predicted fatigue.
-However, fatigue was a very weak indicator of other unmentioned symptoms, despite being explicitly mentioned in 25% of the cases.
-We identified three pairs of symptoms with strong mutual relationships, where the explicit presence of one symptom influenced the implicit estimation of the other, and vice versa.
-These pairs include anhedonia and depressed mood, fatigue and psychomotor agitation/retardation, and anhedonia and concentration difficulties.
-Uncovering GPT-4’s language of depression: GPT-4 generated reliable reasons for diagnosis
-From the left bar chart in Figure 4, it is conclusive that GPT-4 identified explicit markers (positive Cohen’s $d$) of depressive symptoms included words that have been known to be associated with depression from prior studies [46, 47], such as affective words (depressed, feeling low, sad, down), verbs (do), and somatic descriptors (energy, tired, sleep).
-On the other hand, words from non-spans (negative Cohen’s $d$) mostly consisted of functional words and a few nouns (e.g.
-pandemic, home).
-All these words on the left bar chart of Figure 4 had a statistically significant association ($p<0.05$; tested for multiple hypotheses using Bonferroni correction).
-In the right bar chart, 12 out of the 25 words with positive Cohen’s $d$ had statistically significant Pearson correlation ($r;p<0.05$) with GPT-4’s total estimated PHQ-9.
-Essays containing ‘not been depressed’ and mentions of ‘happy’ showed a strong negative association with the total score, and phrases like ‘have been depressed’, ‘sleep’, and ‘low’ (as in low energy) had a positive association.
-From the set of words with negative Cohen’s $d$, only ‘it’ had a significant but weak association ($r=0.09$) with the total estimated PHQ-9.
-This showed that GPT-4 doesn’t strongly depend on any meaningful linguistic signal outside its cited explicit markers.
-3 Discussion
-Receiving professional mental health support is a multi-step, non-linear process that begins with screening and initial assessment.
-A significant percentage of individuals do not receive timely support due to social barriers and infrastructural limitations within the traditional mental healthcare system [48, 49].
-Though effective mental health treatments exist, most individuals with mental disorders do not receive treatment, in part due to shortages of mental healthcare providers [50, 51].
-Digital technologies like GPT-4 which possess unprecedented human-comparable skills and may soon be able to practically address many such challenges.
-With current consumer costs being around \$30/month [52], models such as GPT-4 may be well-positioned to mitigate well-established inhibitors such as high costs, limited mobility, inadequate insurance coverage, and lack of support for parents [15].
-Additionally, it can overcome infrastructural limitations, including the need for cultural adaptation and the dearth of mental health clinicians.
-However, overcoming these challenges does not guarantee precise assessments due to limitations in sensitivity with current measurement tools.
-To enhance accuracy, these technologies can leverage passive and quasi-passive indicators from daily life [53, 54] to form more precise assessments and aid the delivery of timely care to those in need.
-While studies have focused on LLMs’ generations to identify and measure desirable behaviors for mental health support [55, 56], the current study instead sought to inform how the internal structure of depression is represented within an LLM by analyzing its behavior with respect to symptoms.
-Hence, we tasked GPT-4 with identifying the key symptomatic markers of depression expressed within human essays.
-These essays contained exhibits of causes (e.g.
-substance use, sedentary lifestyle) and effects (e.g.
-tired, exhausted) of depression through direct or synonymous mentions (e.g.
-“I don’t like doing things anymore” vs “I don’t see the point of getting out of the bed anymore” both suggest lack of interest).
-From GPT-4’s assessments on language describing depression, we learn how it views the condition to manifest in people, akin to how practicing therapists understand conditions based on the way patients describe it [57, 58].
-We evaluated the viability of GPT-4 for assessing depression using two key principles from measurement theory: reliability and validity.
-We found that GPT-4 displays moderate-high levels of concurrent validity and internal consistency (reliability) with depressive severity from a well-established rating scale measure, PHQ-9.
-The overall result of GPT-4’s convergence with self-report was consistent with overall associations of GPT-4 [59] and other LLM[33] with standard depression measures, while the item-level results extend convergence to detection of depression symptoms.
-Further, we found that GPT-4’s generated explanations were generally faithful [60], with little evidence that it relied on linguistic cues outside its explanations of explicit phrases used to estimate PHQ-9 scores.
-These findings indicate that GPT-4 can be useful at screening patients and identifying symptoms from their language.
-Language is an information-rich medium compared to rating scales [61] which can provide avenues for making hypotheses about predisposing, precipitating, and perpetuating factors of psychopathology for a given patient while estimating the symptom severity.
-Thus, the explanatory output of LLM assessment may free up clinicians to prioritize complex case conceptualization and the creation of carefully tailored treatment plans.
-[…]
-Highly depression-specific symptoms such as anhedonia, depressed mood, and feelings of worthlessness/guilt [22, 62] had both higher base rates of explicit mentions (Table 3) and higher effects on implicit symptoms (Figure 3), explaining the severity estimations when symptoms were unmentioned in the essays.
-Symptoms with lower levels of specificity to depression [22], like sleep and appetite, were inferrable in the presence of almost any other symptom.
-These nonspecific symptoms are shared by depression and anxiety-related disorders [22].
-Another important consideration in explaining LLMs is the instructions it is given.
-Given that the information present in the instructions can affect the performance of these models [63, 64, 65, 66], we instructed GPT-4 to estimate symptoms of depression before estimating the severity of overall depression.
-This provided a means to understand the cohesiveness of the structure of depression, its faithfulness to what it identifies as explicit markers, and varying difficulty levels in estimating the symptoms.
-This mirrors studies of LLMs in other domains that have demonstrated improved effectiveness when allowed to break a task into compositional steps [67, 68] or when instructed to carry out specific modules that culminate into a task [69, 70].
-Using measurement theory to arrive at explanatory schemas can inform how LLMs function, where they are bound to fail, and how to effectively situate them in the mental health support system to maximize overall benefits.
-There has been increased advocacy for transparency and explainability [71] of LLMs to improve public faith and understand the risks associated with its usage, especially in areas such as mental health support [72].
-Our evaluation framework is intended to help clinicians and decision-makers gain insights about model capabilities and pitfalls so that they can determine the capacity at which contemporary models can be used in the current care pipeline and further improve future systems to yield stronger outcomes.
-Our work had several limitations.
-Foremost, the scope of our results is limited to explaining GPT-4’s behaviors when used as a depression assessment tool.
-Although other modalities of behavior beyond the patient’s verbal behavior can be used for the assessment, we employ language because it intrinsically contains more information, and the advent of LLMs enables us to leverage the power of language to deliver accurate assessments [61, 73] Further exploration is required to assess whether these models are suitable for use in other stages of the mental health treatment pipeline.
-It is a multi-stage process, encompassing assessment of several other conditions from time to time, and folding in interventions targeted at addressing the causes of patient’s issues.
-While the nature of the collected data only allowed us to perform a cross-sectional analysis on a small population predominantly from the United Kingdom, we should also evaluate how these models assess mental health for other random samples of the population.
-Accurate mental health assessments would require taking longitudinal symptom trajectories into account as well [35] to overcome the limitations of cross-sectional evaluations, which may fail to capture the dynamic nature of symptom progression and individual variability over time.
-Although transformer-based models have been effective in encoding longitudinal representations of behavioral  [74, 75] and heath markers [76], it is vital to understand how LLMs can effectively disentangle both these trajectories over time to assess the mental health severity.
-Altogether, our analysis of GPT-4 in assessing depressive symptoms highlights two key insights: (a) moderate to high convergence with self-reported data and (b) an underlying symptom schema that explained its assessments.
-GPT-4 ’s schema closely matched self-report for major symptoms, but notably downplayed the link with suicidal ideation and overemphasized psychomotor agitation/retardation.
-This divergence was somewhat mitigated when suicidal ideation was explicitly mentioned, yet the strong overemphasis on psychomotor symptoms led to weaker alignment with both self-reports and expert ratings.
-Our findings indicate that GPT-4 has the potential to assist clinicians by efficiently assessing patients and prioritizing cases based on depression symptom severity, optimizing time for those in urgent need.
-Expanding these methods to other mental health conditions and large language models will deepen our understanding of the benefits and limitations of these tools in enhancing mental health care.
+We applied differential language analysis [40, DLA] to identify the linguistic cues associated with GPT-4’s identification of depressive risk factors.
+Following this, we evaluated whether GPT-4 shows evidence that it uses the phrases it cited as explanations for producing its depression scores.
+Lastly, we used multivariate linear regression to assess how GPT-4 integrates explicitly mentioned factors in essays to estimate risks for unmentioned factors, to uncover its mechanism behind symptom estimation.
+We further apply this framework to GPT-5 and show that performance differences relative to GPT-4 on depression assessment are accounted for by shifts in the learned schema, linking aggregate assessment scores to interpretable internal relations among symptoms.
+Figure 1: Overview of Machine Behavior Evaluation to uncover the structure of depression in GPT-4.
+(1) GPT-4 is prompted with step-by-step instructions on performing depression assessment on human-written essays.
+(2) GPT-4 carried out assessments on data from Gu et al., 2024 [16], containing depression essays written by individuals.
+The model estimated the severity scores for PHQ-9 symptoms explicitly mentioned in the essay, followed by scoring the remaining symptoms that were implicit in the essay.
+(3) GPT-4’s estimated scores and spans are used to infer its internal structure of depressive symptoms, which revealed how the web of symptoms was connected, and what language GPT-4 reliably used to estimate depression severity.
 2 Methods
-Data
+2.1 Data
 Collection.
-Our data was collected as part of another study [17].
-From 1040 participants in total, 258 (24.8%) screened participants reported having been diagnosed with Major Depressive Disorder (MDD) and 259 (24.8%) with Generalized Anxiety Disorder (GAD); and 491 (47.2%) participants were recruited from an unscreened population.
-32 (3.2%) respondents were excluded for not answering the control items (see Measures) correctly.
+Data comes from study participants that were recruited through Prolific[34], an online platform.
+From 1040 participants in total, 258 (24.8%) screened participants reported having been diagnosed with Major Depressive Disorder (MDD), 259 (24.8%) with Generalized Anxiety Disorder (GAD), and 491 (47.2%) participants were recruited from an unscreened population.
+32 (3.2%) respondents were excluded for not answering the control items correctly.
 Removing participants that failed attention check and incomplete survey responses, the dataset comprised 963 participants, 571 identified as female, 388 as male, 2 preferred not to say, and 2 did not respond.
@@
 (n = 679).
-Table S1 shows the summary statistics of the self-reported score for each PHQ-9 symptom in our data.
 The Swedish National Ethics Review Board deemed the research study (Ethics Application 2020-00730) exempt from requiring ethical approval according to Swedish Law (see §§ 3-4 of the Act [2003:460] on ethical review of research involving humans in Sweden).
 The research was carried out according to the Declaration of Helsinki.
-Participants were recruited from the online platform Prolific [77].
 The sample comprised a mixture of screened and unscreened respondents.
+More details on data can be found in Gu et al.,, 2024.
 Measures.
 Below we only describe the measures used in this study – for all details regarding the data please refer to Gu et al., 2023 [16].
-Open-ended question on depression - questions asking participants to describe their depression was adapted from previous research[78] and validated in Gu et al., 2023 [17] (under review).
-The questions read:
+Participants were asked to describe their depression using open-ended language response.
+The questions was adapted from previous research[23] and validated in Gu et al., 2024 [16].
+Responses for the following prompt were used in this work.
 Over the last 2 weeks, have you been depressed or not?
@@
 The rating scale included nine items targeting symptoms such as “Feeling down, depressed, or hopeless”, referring to experiences from the last two weeks.
-A four-point rating scale is used (0 = Not at all to 3 = Nearly every day) to indicate the experienced severity of each symptom.
+A four-point rating scale was used (0 = Not at all to 3 = Nearly every day) to indicate the experienced severity of each symptom.
 Procedure.
@@
 The median time to complete the study was 20.1 minutes (mean=24.0, SD=12.8).
+2.2 Experiment Design
+Prompt.
+Text response describing depression was sent to GPT-4 111versioned 11/06 and referred to as gpt4-1106 in OpenAI’s GPT directory models along with the instructions to carry out assessment (made available as supplementary materials).
+The instruction was composed into three parts:
+  1. 1.
+    Task description: A short paragraph describing the task of estimating the PHQ-9 item scores by analyzing the human-written text.
+The description included the commonly abbreviated form of the 9 items from the PHQ-9 questionnaire.
+  2. 2.
+    Task Steps: A detailed step-by-step instructions dictating GPT-4 to first estimate the scores for the items that are explicitly stated in the text, followed by scoring the items that were not explicitly stated and finally, concluding with the combined PHQ9 score and severity category.
+  3. 3.
+    Output Format: GPT-4 was instructed to format its generation into a predefined data structure (JSON object) to facilitate easy processing of data for analysis.
+Although, GPT-4 generated outputs fulfilled the completion of PHQ-9 estimation for all the data, it was not completely faithful to the details of the instructions.
+There were a few issues observed with following the steps detailed in the instructions, especially step 1.
+59 (6.1%) out of 956 samples, had small issues that were manually rectified.
+These included cases where GPT-4 had
+  1. 1.
+    made subtle changes to the symptom abbreviations (E.g. ’Feelings of Worthlessness or Guilt’ instead of ’Worthlessness or Guilt’),
+  2. 2.
+    identified specific aspects of the symptoms (E.g.
+’Insomnia’ in ’Insomnia or Hypersomnia’, ’Psychomotor Retardation’ in ’Psychomotor Agitation/Retardation’, ’Poor Appetite’ in ’Poor Appetite or overeating’ etc.),
+  3. 3.
+    split the individual steps into separate JSONs, instead of nesting them within a single JSON as described in the prompt.
+  4. 4.
+    included a summary of the assessment which was a short essay of its JSON output
+  5. 5.
+    included symptoms that were not a part of the PHQ-9 (E.g. Financial Worries, Social Exclusion)
+For the first two cases, the symptoms were manually corrected to match the abbreviations in Step 1 of the instructions.
+Type 3 generations were rectified by manually reformatting, and the summaries in type 4 were removed.
+The symptoms in type 5 generations were left out of the analysis.
+GPT-4’s generations varied based on its hyperparameters.
+For this work, the temperature value was set to 0.0 to select the most likely token at each step and to allow reproducibility of experiments.
+Other hyperparameters include frequency penalty (0.1), presence penalty (0.0), top p (1.0), and max tokens (450).
+Cleaning and Filtering GPT-4’s Responses.
+GPT-4 generated responses in JSON format (example output available in supplementary materials), were turned into tabular form by extracting all the scores into individual columns along with a binary variable indicating whether a symptom was found to be explicit by GPT-4.
+The excerpts from the user-written text identified by GPT-4 as explicit symptom markers were extracted using simple regular expressions.
+GPT-5 also used the same experimental design and processing steps.
+Expert Judgments.
+Two psychology experts independently assessed the depression severity from the depression texts on a random subsample of 209 instances.
+They were instructed to read the essay to identify excerpts indicative of PHQ-9 symptoms and estimate a severity score for them.
+If there were no indicative excerpts, they were asked to estimate the symptom severity based on the essay and the estimated explicit symptom scores.
+The symptoms accompanied by excerpts would pertain to explicit markers as identified by experts.
+2.3 Analytic Plan
+Our analytical approach draws from measurement theory and behavioral analysis to systematically understand GPT-4’s and GPT-5’s representation of depressive symptoms.
+We used three strategies to bring out different aspects of how the models processes and integrates depressive symptoms from natural language.
+We begin by assessing the convergent validity between GPT-4 and GPT-5 estimates, self-reported PHQ-9 scores, and expert judgments using Pearson correlations to establish the basic reliability of language models’ assessments before examining its underlying mechanisms.
+Following this, the rest of the analysis in the main paper focuses on GPT-4, and GPT-5’s analysis can be found in the Supplementary section.
+We perform latent structure analysis to model how symptoms are related within GPT-4.
+Subsequently, we use Differential Language Analysis to gain a deeper understanding of how it views depression in language, and forms severity assessments.
+Finally, we model the process behind the estimation of implicit symptoms’ severity based on explicit symptom estimates.
+2.3.1 Latent Structure Analysis
+Inter-symptom Relationships.
+To understand whether GPT-4’s assessment behavior reflects symptom associations consistent with established depression research, we computed Pearson correlations between all pairs of the nine PHQ-9 symptoms for both GPT-4 estimates and self-reported scores.
+This reveals how GPT-4 conceptualizes the relationships among different depressive symptoms compared to how individuals experience and report these relationships.
+The correlation matrices allow us to identify whether GPT-4 has learned symptom associations that align with or deviate from established psychometric patterns in depression research.
+To contrast the inter-symptom relationships between GPT-4 and self-report, we computed the difference between the inter-symptoms relationship matrices.
+For statistical robustness, this difference matrix was computed over 500 trials by bootstrapped resampling to extract the 95% confidence interval of the differences.
+CIs that included 0 were regarded statistically insignificant difference in relationship between the two symptoms $(p>.95)$ across GPT-4 and self-report.
+Differential Item Functioning Analysis.
+We conducted Differential Item Functioning (DIF) analysis using Item Response Theory to compare how effectively each symptom discriminates between varying levels of depression severity across GPT-4 estimates versus self-reported scores.
+In IRT, data from ordinal items are modeled using latent variables with logistic link functions [49].
+For this analysis, we used the 1-parameter logistic IRT model (Rasch model) [36] which constrains the slope parameter to 1.0, with item location parameters estimated from responses.
+The item location parameter indicates where each symptom functions most effectively along the latent depression continuum.
+DIF analysis examines whether items function differently across groups by comparing rank orderings of item locations.
+Lower ranks indicate symptoms that discriminate better at lower depression severity levels, while higher ranks indicate symptoms most informative for distinguishing individuals with severe depression.
+We computed DIF ranks for both GPT-4 and self-report assessments, with rank differences revealing whether GPT-4’s language-based assessment emphasizes different symptoms than traditional self-report measures across the depression severity spectrum.
+To ensure statistical robustness, we estimated 95% confidence intervals around rank differences using bootstrapped resampling over 500 trials, allowing us to identify statistically significant differences in how symptoms function across assessment methods.
+2.3.2 Explicit Symptom Identification and Linguistic Analysis
+Language Analysis of Depression Markers.
+Beyond examining GPT-4’s numerical estimates, we investigated whether the model’s scoring behavior aligns with linguistic patterns known to characterize depression in human language, particularly focusing on language features that extend beyond GPT-4’s explicitly identified symptom markers.
+For each participant’s text, we computed 1-, 2-, and 3-gram relative frequencies, categorizing tokens into two groups: those present within GPT-4’s explicit spans (text segments the model cited as evidence) and those outside these spans.
+To ensure robust analysis, we filtered out N-grams used by fewer than 10 participants, and removed any non alphabetic characters (including numbers or special characters).
+We then computed Cohen’s d [6] effect sizes comparing n-gram relative frequencies within explicit spans versus the remainder of user text.
+This analysis identifies linguistic features that GPT-4 systematically attributes as an explicit marker of depression.
+To control for multiple comparisons, we applied Benjamini-Hochberg correction, which maintains acceptable false discovery rates when testing numerous linguistic features simultaneously.
+We further refined the analysis by filtering features using Pointwise Mutual Information [5] with a threshold of 4.0, which identifies n-grams with strong associative relationships to explicit symptom markers while removing spurious associations.
+Finally, we correlated the relative frequencies of these linguistically significant features with GPT-4’s total PHQ-9 scores to determine whether the model’s severity estimates rely on language patterns beyond its explicitly cited evidence.
+2.3.3 Mechanism Analysis: Explicit-Implicit Symptom Interplay
+Multivariate Regression Modeling of Symptom Inference.
+To understand GPT-4’s mechanism for estimating symptoms that are not explicitly mentioned in user text, we modeled how the GPT-4’s implicit symptom scores relate to its explicit symptom identifications.
+For each symptom $i$ from the complete set of PHQ-9 symptoms $P$, we fitted a separate multivariate linear regression model predicting GPT-4’s estimated score for symptom $i$, specifically in cases where that symptom was not explicitly identified by the model.
+The regression models used scores of all other explicitly identified symptoms as predictors, creating a comprehensive picture of how GPT-4 integrates explicit information to infer unmentioned symptoms.
+When a predictor symptom was not explicitly identified, we replaced its score with the dataset mean for that symptom.
+$\beta$ represent the strength and direction of association between each explicit symptom and the implicit estimation of the target symptom:
+$s_{i}^{imp}=\sum_{\begin{subarray}{c}j\in P\\ j\neq i\end{subarray}}\beta_{j}s_{j}^{exp}$ (1)
+This approach reveals GPT-4’s learned patterns about symptom co-occurrence, similar to how a clinician might infer risk for sleep problems when someone explicitly reports severe mood symptoms.
+To ensure statistical robustness and account for sampling variability, we repeated this entire modeling process 500 times using bootstrapped samples of the data.
+Coefficients whose 90% confidence intervals included zero were considered statistically non-significant and excluded from interpretation, providing conservative estimates of GPT-4’s symptom inference mechanisms.
+3 Results
+GPT-4 analyzed language samples describing depressive feelings from 955 individuals to estimate symptom-level scores based on the PHQ-9 questionnaire [27].
+First, the model identified explicitly described symptoms, citing relevant text spans from the language sample and assigning severity scores.
+For the remaining PHQ-9 symptoms that were not explicitly mentioned, GPT-4 was instructed to provide rationales along with its inferred severity estimates.
+Participants independently completed the PHQ-9 questionnaire reflecting symptom severity over the previous two weeks.
+Additionally, two psychology experts manually annotated symptoms in a random subsample of 209 responses following the same guidelines given to GPT-4.
+GPT-4 estimates were compared against self-reported scores on the full dataset and against expert annotations in the subsample (Table 1).
+Convergent Validity: GPT-4 estimates correlate with experts-judgments and self-report.
+Table 1 shows the average Pearson correlation ($r$) between GPT-4 estimates and experts for each symptom, as well as the correlation between GPT-4 and self-reported scores, both computed on the subset assessed by the experts ($N=209$).
+GPT-4 demonstrated high agreement with self-report ($r=0.70$) and experts-judgments (avg $r=0.81$) on total PHQ-9 ratings.
+Across all symptoms, the assessors of depression essays (i.e., GPT-4 and the experts) demonstrated higher agreement with each other than either did with self-reported PHQ-9 scores.
+The absolute difference between GPT-4’s agreements with self-report and experts (difference between first two columns) was less than $0.10$ for all somatic symptoms (fatigue, appetite, and psychomotor agitation/retardation), except for issues related to Sleep ($\Delta=0.15$).
+All cognitive and emotional symptoms except for anhedonia had a difference in correlation scores over 10 Pearson points.
+Highest differences were seen with suicidal ideation ($\Delta=0.32$), worthlessness/guilt ($\Delta=0.20$) and concentration difficulties ($\Delta=0.14$).
+GPT-4’s agreement with experts ($r_{exp1}=0.80,r_{exp2}=0.81$) on total PHQ-9 score was very similar to the agreement between the experts ($r=0.83$).
+GPT-5 showed consistently lower agreement with self-report than GPT-4 across all symptoms and overall PHQ-9 score.
+Symptom-Symptom Relationships: GPT-4 and Self-report have similar inter-symptom covariates overall, except for two symptoms.
+The symptom-symptom associations from GPT-4 estimates in 2(a) (left) had positive Pearson correlations across the board (r=0.230.78r=0.23-0.78).
+This was in agreement with the symptom-symptom associations from self-reported (right) scores (r=0.380.78r=0.38-0.78).
+The effect sizes of the symptom-symptom relationships from GPT-4 estimates were similar to self-report for all but two symptoms, namely, psychomotor agitation/retardation and suicidal ideation.
+2(b) clearly shows this difference between symptom-symptom associations from GPT-4 and self-report is statistically significant (95% CI doesn’t include 0.0).
+GPT-4-estimated severities of psychomotor agitation/retardation had a stronger association with other depressive symptoms than self-report.
+In contrast, suicidal ideation was less related to other depressive symptoms for GPT-4 than in the self-report.
+This is evident from the significantly lower association of suicidal ideation with the rest of the symptoms than what was observed in self-reported scores.
+Differential Item Functioning Analysis: GPT-4 and Self-report have similar discriminant characteristics over latent depression scores.
+We conducted Differential Item Functioning (DIF) analysis using Item Response Theory (IRT) to compare discriminative power between self-report and GPT-4-derived symptom scores [47, 38, 2].
+DIF ranks reflect how effectively each symptom distinguishes between varying levels of latent depression severity: lower ranks indicate better discrimination at lower severity, and higher ranks indicate better discrimination at higher severity.
+A positive rank difference (GPT-4 minus self-report) implies that GPT-4’s language-based assessment discriminates better at higher depression severity levels compared to self-report, while a negative difference indicates the opposite.
+DIF ranks for both methods, along with their 95% confidence intervals estimated via bootstrapped resampling over 500 trials, are presented in 2(c).
+Overall, the discriminative characteristics of items were consistent between GPT-4 and self-report assessments.
+Suicidal ideation, appetite, and psychomotor agitation/retardation had higher ranks (7–9), indicating these symptoms were most informative in distinguishing individuals with higher latent depression severity.
+On the other hand, anhedonia, depressed mood, and fatigue had lower ranks (1–3), signifying these symptoms effectively discriminated between individuals experiencing milder or lower levels of depression.
+All item ranks showed narrow 95% confidence intervals, reflecting high reliability.
+Suicidal ideation (rank = 9), sleep (rank = 4), and anhedonia (rank = 3) exhibited identical ranks between GPT-4 and self-report.
+Rank differences for other symptoms were minor, reinforcing that language-based severity scores from GPT-4 closely align with the discriminative item characteristics observed in self-reported PHQ-9 scores.
+Explicit vs. Implicit Symptoms: GPT-4 Estimates Explicitly Mentioned Symptoms More Precisely.
+Agreement between GPT-4 estimates and self-report was consistently higher when symptoms were explicitly mentioned, showing an average increase of +0.18+0.18 Pearson correlation points over implicitly inferred symptoms –— even though explicit mentions were identified in only about one-quarter of instances (3(a)).
+Notably, explicit mentions significantly improved GPT-4’s accuracy ($\geq 0.25$ Pearson points) in estimating anhedonia, depressed mood, and suicidal ideation compared to implicit inferences.
+Language explicitly describing these three symptoms enabled more precise estimates from GPT-4, despite substantial variation in how frequently they were identified to be explicitly mentioned (range: 5.55%–96.44%).
+The cardinal symptoms, depressed mood and anhedonia, had particularly high rates of explicit mentions (96% and 53%, respectively), considerably surpassing other symptoms.
+Uncovering GPT-4’s language of depression: GPT-4’s cited evidences were reliable
+The left bar chart of 3(b) shows that words identified by GPT-4 as explicit markers of symptoms (positive Cohen’s $d$ values).
+These include emotion words (e.g., ”depressed,” ”sad,” ”down”), verbs (e.g., ”do”), and somatic descriptors (e.g., ”tired,” ”sleep,” ”energy”).
+In contrast, words that GPT-4 rarely cited as explicit markers (negative Cohen’s $d$) were mainly neutral functional words and nouns (e.g., ”pandemic,” ”home”).
+All these highlighted words showed statistically significant associations after correcting for multiple comparisons (Benjamini-Hochberg corrected, $p<0.05$).
+The bar chart on the right further shows that 10 of the 17 explicitly cited words significantly correlated with GPT-4’s estimated total PHQ-9 scores ($p<0.05$).
+Phrases such as ”have been depressed,” ”sleep,” and ”low” (in ”low energy”) positively correlated with higher severity scores, while phrases like ”not been depressed” and ”happy” negatively correlated, indicating lower severity.
+Among words not explicitly cited, only ”it” showed a small but significant correlation ($r=0.09$), suggesting GPT-4 primarily bases its severity scores on explicitly cited meaningful linguistic markers.
+Interplay between explicitly mentioned symptoms and inferred symptoms.
+We examined how explicit symptoms influence implicit symptoms in GPT-4’s assessments by calculating the regression coefficients ($\beta$) from a multivariate linear regression model.
+This linear model predicted implicit symptom scores based on explicit symptom scores while controlling for all other implicit symptoms.
+Figure 4 depicts the effect sizes of explicit symptoms in influencing GPT-4’s estimate of implicit symptoms.
+When depressed mood was not explicitly mentioned, it was largely inferred from explicit mentions of anhedonia ($\beta=.60$).
+Conversely, anhedonia was influenced by depressed mood ($\beta=.49$) and concentration difficulties ($\beta=.24$).
+The cardinal symptoms of depression were common explicit drivers of nearly all other symptoms with large effect sizes, underscoring their central role in the network of depressive symptoms.
+This shows that depressed mood and anhedonia are pivotal for accurately estimating the severity of other less specific symptoms of depression, such as sleep disturbances, appetite changes and fatigue, particularly when they are not directly mentioned in the text.
+Furthermore, when individuals mentioned sleep disturbances, GPT-4 moderately used this to estimate other somatic symptoms, such as fatigue ($\beta=.18$) and abnormal appetite ($\beta=.24$).
+However, the reverse relationship wasn’t significant: sleep disturbances were not inferred from the other somatic markers.
+However, other physical symptoms didn’t strongly predict sleep disturbances.
+Instead, sleep issues were primarily inferred from depressed mood ($\beta=.26$).
+All symptoms except for fatigue that were explicitly mentioned more than 10% of the time —– including the cardinal symptoms and feelings of worthlessness/guilt —– significantly helped GPT-4 to estimate the severity of unmentioned symptoms.
+The explicit mention of nearly all the items predicted fatigue.
+However, fatigue was a very weak indicator of other unmentioned symptoms, despite being explicitly mentioned in 1 out of 4 cases.
+This reflects fatigue’s nonspecific nature, as it occurs broadly across many conditions and isn’t uniquely informative for depression severity.
+Three pairs of symptoms showed strong two-way relationships, each explicitly mentioned symptom significantly influencing the other’s implicit estimation and vice versa: anhedonia and depressed mood, fatigue and psychomotor agitation/retardation, and anhedonia and concentration difficulties.
+4 Discussion
+Receiving professional mental health support is a multi-step, non-linear process that begins with screening and initial assessment.
+A significant percentage of individuals do not receive timely support due to social barriers and infrastructural limitations within the traditional mental healthcare system [7, 1].
+Though effective mental health treatments exist, most individuals with mental disorders do not receive treatment, in part due to shortages of mental healthcare providers [20, 25].
+Digital technologies like GPT-4 which possess unprecedented human-comparable skills and may soon be able to practically address many such challenges.
+With current consumer costs being around \$20/month, models such as GPT-4 may be well-positioned to mitigate well-established inhibitors such as high costs, limited mobility, inadequate insurance coverage, and lack of support for parents [45].
+Additionally, it can overcome infrastructural limitations, including the need for cultural adaptation and the dearth of mental health clinicians.
+While studies have focused on identifying and measuring desirable behaviors of LLMs for mental health support [4], the current study instead sought to inform how the internal structure of depression is represented within an LLM by analyzing its behavior with respect to symptoms.
+Hence, we tasked GPT-4 with identifying the key symptomatic markers of depression expressed within human essays.
+These essays contained exhibits of causes (e.g.
+substance use, sedentary lifestyle) and effects (e.g.
+tired, exhausted) of depression through direct or synonymous mentions (e.g.
+“I don’t like doing things anymore” vs “I don’t see the point of getting out of the bed anymore” both suggest lack of interest).
+From GPT-4’s assessments on language describing depression, we learn how it views the condition to manifest in people, akin to how practicing therapists understand conditions based on the way patients describe it [10, 44].
+We examined the feasibility of using GPT-4 for depression assessment by evaluating two fundamental psychometric properties: reliability and validity.
+GPT-4 demonstrated strong concurrent validity, indicating its assessments align closely with established measures of depression, such as the PHQ-9 questionnaire.
+Additionally, GPT-4 showed high internal consistency, suggesting it reliably captures the interconnected nature of depressive symptoms.
+Our findings align with previous research showing that LLMs [58, 14] generally correlate well with standard depression scales, and we extend these results to symptom-level assessments.
+Importantly, GPT-4’s cited explicit markers demonstrated both validity and reliability: the linguistic markers it identified aligned with prior findings [9, 46], and were positively associated with its severity estimates, whereas uncited language from depression essays had no meaningful associations.
+Such language-based assessments offer richer and more detailed insights than traditional rating scales [24], which provide avenues for making hypotheses about predisposing, precipitating, and perpetuating factors of psychopathology enabling a more personalized treatment planning.
+Going beyond evaluating GPT-4’s convergence with self-reports and experts, we inferred its schema of symptoms from its behavior in assessing depression.
+This schema, a representation of the psychological constructs, that was derived from GPT-4’s responses, was depicted as a web of associations between symptoms.
+We found GPT-4 and self-report shared commonalities in their schema across the first seven PHQ-9 symptoms, while they differed for psychomotor agitation/ retardation and suicidal thoughts.
+Specifically, GPT-4 ’s assessment of suicidal thoughts closely matched expert evaluations and explicit self-reports but struggled to infer it when not directly mentioned.
+This difficulty highlights the challenge of inferring severe but less frequently mentioned symptoms solely from associated emotional states like depressed mood or anhedonia.
+In contrast, GPT-4 consistently showed lower convergence in assessing psychomotor disturbances, suggesting that these symptoms are less observable through language compared to others.
+This finding aligns with prior research indicating that objective behavioral signals, such as speech patterns or facial expressions, are more reliable for assessing psychomotor symptoms due to potential recall biases in self-reports [43].
+Notably, both suicidal ideation and psychomotor disturbances demonstrated high discriminant power with respect to latent depression severity (2(c)), underscoring their clinical importance and the need to better understand how GPT-4 diverges in modeling these constructs.
+GPT-4 ’s approach to inferring symptoms that are not explicitly mentioned aligns with prior research showing that some symptoms are more central and uniquely tied to depression, while others are broader and commonly appear across different mental health conditions [55].
+Highly depression-specific symptoms such as anhedonia, depressed mood, and feelings of worthlessness/guilt [55, 17] were frequently mentioned in language (3(a)) and strongly influenced GPT-4’s inference on other symptoms (Figure 4).
+This pattern suggests these core symptoms serve as foundational indicators in GPT-4 ’s schema of depression, mirroring clinical practice where these cardinal symptoms heavily guide diagnostic reasoning [11].
+Conversely, nonspecific symptoms, such as sleep disturbances and appetite changes – common to both depression and anxiety disorders – could be reliably inferred by GPT-4 from the presence of almost any explicitly mentioned symptom.
+Another important consideration in explaining LLMs is the instructions it was given.
+As the information in instructions can affect the model’s response [3], we instructed GPT-4 to estimate symptoms of depression before the overall depression severity.
+This provided a means to understand the cohesiveness of the structure of depression, its faithfulness to what it identifies as explicit markers, and varying difficulty levels in estimating the symptoms.
+This mirrors studies of LLMs in other domains that have demonstrated improved effectiveness when allowed to break a task into compositional steps [56] or when instructed to carry out specific modules that culminate into a task [22].
+There has been increased advocacy for transparency and explainability [57] to improve public faith and understand the risks it comes with, especially in areas such as mental health support [29].
+Using measurement theory to arrive at explanatory schemas can inform how LLMs function, where they are bound to fail, and how to effectively situate them in the mental health support to maximize overall benefits.
+Our evaluation framework is intended to help clinicians and decision-makers gain insights about model capabilities and pitfalls to determine the capacity at which contemporary models can be used in the care pipeline and further improve future systems to yield stronger outcomes.
+Our work had several limitations.
+Foremost, the scope of our results is limited to explaining GPT-4’s behaviors when depression is conceptualized through the lens of the PHQ-9.
+While the PHQ-9 is a widely used and validated tool, it reflects a categorical, DSM-based view of depression that may not fully capture the complexity of psychological distress.
+Future work can expand this approach to data-driven, dimensional models such as HiTOP [26], which may better capture the heterogeneity and complexity of psychopathology across individuals.
+Although other modalities beyond language (e.g., speech, behavior) can enrich assessment, we focus on language due to its information-rich nature and the ability of LLMs to leverage this medium for making accurate assessments [24, 52].
+Additionally, our study was limited to a cross-sectional analysis of a relatively small, predominantly UK-based sample.
+Broader validation across diverse populations and contexts is necessary.
+Accurate mental health assessments would require taking longitudinal symptom trajectories into account [15] to overcome the limitations of cross-sectional evaluations, which fails to capture the dynamic nature of symptom progression and individual variability over time.
+Although transformer-based models have been effective in encoding longitudinal representations of behavioral [59] and health markers [51], it is vital to understand how effective are LLMs in disentangling both these trajectories over time to assess the mental health.
+5 Conclusion
+This study uncovered the underlying symptom schema within GPT-4 that explains how its arrived at scores having high convergence with experts and standard self-report tests.
+While GPT-4 ’s schema closely matched self-report for major symptoms, it notably downplayed the link with suicidal ideation and overemphasized psychomotor agitation/retardation.
+This divergence was somewhat mitigated when suicidal ideation was explicitly mentioned, suggesting it is more conservative when inferring symptoms of suicidality without explicit evidence.
+On the other hand, its overemphasis on psychomotor symptoms seemed to blanket bias, happening whether it identified evidence or not.
+With awareness of these discrepancies, the results suggest GPT-4 has the potential to assist clinicians, such as in screening to optimize time for those in urgent need.
+More widely, this work demonstrates how a machine behavior approach yields explanations of large language models’ processing of psychological concepts.
+Ethics Approval and Consent to participate
+This study was deemed exempt from IRB approval. All study participants’ data was anonymized and they consented to share the data for research.
+Declarations
+Funding.
+AVG and RLB were supported in part by NIH grant R01-AA028032, VV was supported in part by a DARPA grant #W911NF-20-1-0306, and HAS was supported in part by both the grants.
+YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290.
+JCE was supported by Stanford’s Institute for Human-Centered AI.
+LF was supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence research on LLMs, and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University.
+ONEK, KK and VCE were supported by FORTE (STY-2022/0007; 2022-01022).
+The conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, NIH, any other government organization, or the U.S./ German Government.
+CoI/Competing Interest. ONEK and KK co-founded a start-up that uses computational language assessments to diagnose mental health problems.
+Preprints
+A preprint of this article is available online: arXiv.2411.13800
+Data availability
+The data for this project will be made available through Github.
+All data including the original study data, GPT-4 responses, cleaned version, as well as the processed features will be open-sourced for replication.
+All programs for this work were coded in Python and will be open-sourced with the data in the same Github repository.
+This includes programs that accessed GPT-4 to generate responses using OpenAI’s API, cleaning the data, and our analysis.
+Acknowledgments
+The authors extend their gratitude to Daniel Klein, Roman Kotov, Nicholas Eaton, Dina Vivian, Whitney Ringwald, August Nilsson, and Siddharth Manglik for providing feedback on this work.
+Supplementary
+The distribution of severity scores for self-report and GPT-4’s assessments over 955 samples from Gu et al., 2024 [16] is shown in Table 2.
+While, GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms, showing a different distribution at higher values.
+This shows that PHQ-9 displays an under-estimation bias of severity scores.
+Table 2: Distribution of self-report and GPT-4 estimated scores for each item on PHQ-9 (N=955).
+GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, but its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms.
+Item % Scoring 0 % Scoring 1 % Scoring 2 % Scoring 3
+Self Report GPT-4 Self Report GPT-4 Self Report GPT-4 Self Report GPT-4
+Anhedonia 24.3 28.5 32.5 43.4 24.4 26.5 18.8 1.7
+Depressed Mood 23.8 25.0 32.1 24.6 22.0 44.0 22.2 6.5
+Sleep 25.1 35.2 23.8 54.4 21.1 10.1 30.0 0.4
+Fatigue 15.4 29.3 27.2 50.9 22.7 19.4 34.6 0.4
+Appetite 30.6 57.2 25.9 40.0 19.1 2.7 24.4 0.1
+Worthlessness / Guilt 29.4 43.3 25.4 37.5 23.0 18.3 22.2 0.9
+Difficulty concentrating 28.0 36.0 30.9 58.3 24.1 5.7 16.9 0.0
+Psychomotor Agi./ Ret. 55.5 53.3 23.2 45.3 15.0 1.4 6.4 0.0
+Suicidal ideation 60.3 87.9 21.4 8.8 10.0 2.3 8.3 1.1
+Figure 5: Prompt given to the study Participants to express their feeling of depression using language.
+Over the last 2 weeks, have you been depressed or not?
+Please answer the question by typing at least a paragraph below that indicates whether you have been depressed or not.
+Try to weigh the strength and the number of aspects that describe if you have been depressed or not so that they reflect your overall personal state of depression.
+For example, if you have been depressed, then write more about aspects describing this, and if you have not been depressed, then write more about aspects describing that.
+Write about those aspects that are most important and meaningful to you.
+Write at least one paragraph in the box.
 Figure 6: Instruction to GPT-4 with an example input text.
@@
 STEP 1: { Depressed Mood: [‘I feel overwhelmed’, 2], Worthlessness or Guilt: [‘annoyed with myself for allowing it to continue’, 2], Difficulty Concentrating: [‘my emotions are unpredictable’, 0] } STEP 2: { Anhedonia: [“Based on the overall tone of hopelessness and resentment, it is likely that the author is experiencing diminished interest or pleasure”, 1], Insomnia or Hypersomnia: [“No explicit mention, but stress and emotional unpredictability may affect sleep patterns”, 1], Fatigue: [“Feeling overwhelmed could imply a state of mental fatigue”, 1], Poor appetite or overeating: [“No explicit mention, but stress often affects eating habits”, 1], Psychomotor Agitation or Retardation: [“‘avoiding talking to people’ may suggest a withdrawal that could be associated with psychomotor retardation”, 1], Suicidal Ideation: [“No explicit mention of suicidal thoughts or self-harm”, 0] } STEP 3: { Final: [Combined Score: 9, Category: Moderate] }
-Experiment Design
-Prompt.
-Only the text response prompted to describe one’s depression was sent to GPT-4 111versioned 11/06 and referred to as gpt4-1106 in openAI’s GPT directory along with the instruction described in Table 5.
-The instruction was composed into three parts:
-  1. (A)
-    Task description: A short paragraph describing the task of estimating the PHQ-9 item scores by analyzing the human-written text.
-The description included the commonly abbreviated form of the 9 items from the PHQ-9 questionnaire.
-  2. (B)
-    Task Steps: A detailed step-by-step instructions dictating GPT-4 to first estimate the scores for the items that are explicitly stated in the text, followed by scoring the items that were not explicitly stated and finally, concluding with the combined PHQ9 score and severity category.
-  3. (C)
-    Output Format: GPT-4 was instructed to format its generation into a predefined data structure (JSON object) to facilitate easy processing of data for analysis.
-Although, GPT-4 generated outputs fulfilled the completion of PHQ-9 estimation for all the data, it was not completely faithful to the details of the instructions.
-There were a few issues observed with following the steps detailed in the instructions, especially step 1.
-59 (6.1%) out of 956 samples, had small issues that were manually rectified.
-These included cases where GPT-4 had
-  1. (i)
-    made subtle changes to the symptom abbreviations (E.g. ’Feelings of Worthlessness or Guilt’ instead of ’Worthlessness or Guilt’),
-  2. (ii)
-    identified specific aspects of the symptoms (E.g.
-’Insomnia’ in ’Insomnia or Hypersomnia’, ’Psychomotor Retardation’ in ’Psychomotor Agitation/Retardation’, ’Poor Appetite’ in ’Poor Appetite or overeating’ etc.),
-  3. (iii)
-    split the individual steps into separate JSONs, instead of nesting them within a single JSON as described in the prompt.
-  4. (iv)
-    included a summary of the assessment which was a short essay of its JSON output
-  5. (v)
-    included symptoms that were not a part of the PHQ-9 (E.g. Financial Worries, Social Exclusion)
-For the first two cases, the symptoms were manually corrected to match the abbreviations provided in Step 1 of the instructions.
-Type iii generations were rectified by manually reformatting the JSON, and the summaries in type iv generations were removed.
-The symptoms in type v generations were left out of the analysis.
-GPT-4’s generations varied based on its hyperparameters.
-For this work, the temperature value was set to 0.0 to select the most likely token at each step and to allow reproducibility of experiments.
-Other hyperparameters include frequency penalty (0.1), presence penalty (0.0), top p (1.0), and max tokens (450).
-Cleaning and Filtering GPT-4’s Responses
-GPT-4 generated responses which were in JSON format as shown in Figure 6 were turned into tabular data by extracting all the scores into individual columns along with a binary variable indicating whether a symptom was found to be explicit by GPT-4.
-Table S1 shows the descriptive statistics of the individual symptom scores estimated by GPT-4.
-The excerpts from the user-written text identified by GPT-4 as explicit symptom markers were extracted using simple regular expressions and stored as a CSV.
-Expert Judgements
-Two psychology experts independently assessed the depression severity from user-written text on a random sample of 209 instances from the same dataset.
-They were instructed to read the essay to identify excerpts indicative of PHQ-9 symptoms and estimate a severity score for them.
-If there were no indicative excerpts, they were asked to estimate the symptom severity based on the essay and the estimated explicit symptom scores.
-The symptoms accompanied by excerpts would pertain to explicit markers as identified by experts.
-Analysis
-Item Response Theory.
-In IRT, data from dichotomous or ordinal items are modeled using one or more latent variables with nonlinear link functions, most commonly logistic functions [79].
-For this analysis, we used the 1-PL dichotomous logistic IRT model, aka Rasch model [80] to fit item parameters and assess the presence of each of the nine symptoms.
-This model constrains the slope parameter of the logistic function to 1.0, with the intercept or item location estimated from the responses.
-The item location parameter indicates where each item stands in relation to assessing depression.
-We analyzed the rankings of item locations to compare the two sets of scores across all 955 participants.
-Interplay Between Explicit and Implicit Symptoms.
-For each symptom $i$ from the set of all symptoms $P$ (as defined by the PHQ-9), we trained a linear regression model to predict the GPT-4-estimated score of symptom $i$ when it was not explicitly identified.
-The model was trained to minimize the sum of squared errors, using the scores of all other explicitly identified symptoms $s_{j}^{exp}$, where $j\in P$ and $j\neq i$, as predictors.
-The linear regression coefficients $\beta$ represent the association between the implicit score of symptom $i$ and the explicit scores of the remaining symptoms, specifically when symptom $i$ was not explicitly identified by GPT-4 in the user text:
-$s_{i}^{imp}=\sum_{\begin{subarray}{c}j\in P\\ j\neq i\end{subarray}}\beta_{j}s_{j}^{exp}$ (1)
-In cases where the score for $s_{j}^{exp}$ was not explicit (i.e., not identified by GPT-4), it was replaced by the mean score for that symptom.
-This process was repeated 500 times using bootstrapped samples of the data.
-Coefficients whose 90% confidence intervals included 0 were excluded from Figure 3.
-Language Of Depression.
-For each user-written text, 1, 2, and 3 grams usage was computed as a relative frequency for each user-written text split into two groups, ie., tokens that were present in the GPT-4’s excerpt spans and tokens that were left out of the spans.
-We removed all n-grams used by less than 10 participants and limited them to n-grams that were only composed of alphabets (filtering n-grams that had numbers and special characters).
-We computed Cohen’s d [81] using n-grams’ relative frequencies when it was found in explicit spans and the rest of the user text.
-After applying Bonferroni correction, it resulted in 56 tokens associated with explicit indication of symptoms (positive d) and 30 that were implicit (negative d) that were statistically significant ($p<0.05$).
-We computed the n-gram (1, 2, and 3) relative frequencies for each user-written text, and created two separate feature sets containing the relative frequencies of positive and negative Cohen’s d respectively.
-Both features were filtered to n-grams with a Pointwise Mutual Information [82] of at least 4.0, which reduced it to 45 and 25 features respectively.
-To uncover if GPT-4 estimates rely on any linguistic signals outside the explicit markers it generated, these sets of positive and negative words’ relative frequencies (computed per user) were correlated with the total PHQ-9 score from GPT-4.
-This analysis was performed using DLATK[83], a Python library for language analysis.
-Declarations
-[…]
-Funding AVG and RLB were supported in part by NIH grant R01-AA028032, VV was supported in part by a DARPA grant #W911NF-20-1-0306, and HAS was supported in part by both the grants.
-YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290.
-JCE was supported by Stanford’s Institute for Human-Centered AI.
-LF was supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence research on LLMs, and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University.
-ONEK, KK and VCE were supported by FORTE (STY-2022/0007; 2022-01022).
-The conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, NIH, any other government organization, or the U.S./ German Government.
-[…]
-CoI/Competing Interest ONEK and KK co-founded a start-up that uses computational language assessments to diagnose mental health problems.
-[…]
-Ethics Approval and Consent to participate This study was deemed exempt from IRB approval. All study participants’ data was anonymized and they consented to share the data for research.
-[…]
-Data availability The data for this project will be made available through Github.
-All data including the original study data, GPT-4 responses, cleaned version, as well as the processed features will be open-sourced for replication.
-[…]
-Materials availability All materials including the manuscript, display Items (tables and figures), code, and data will be available to the public through Github.
-[…]
-Code availability All programs for this work were coded in Python and will be open-sourced with the data in the same Github repository.
-This includes programs that accessed GPT-4 to generate responses using OpenAI’s API, cleaning the data, and our analysis.
-[…]
-Author Contributions AVG, LF, and HAS designed the research; AVG, YKL, and VV built the analytic tools; ONEK and KK were involved in data collection using Prolific; Clinical psychology experts VCE and KK carried out the expert judgements on the depression essays; AVG, VV, YKL, ECS, JCE, KK, ONEK, VCE, HAS, RLB and LF performed the analysis and wrote the manuscript; TD and AVG made all the display items; All authors made the final decision to submit the manuscript.
-Supplementary
-The distribution of severity scores for self-report and GPT-4’s assessments over 955 samples from Gu et al., 2023 [17] is shown in  Table S1.
-While, GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms, showing a different distribution at higher values.
-This shows that PHQ-9 displays an under-estimation bias of severity scores.
-Table S1: Distribution of self-report and GPT-4 estimated scores for each item on PHQ-9 (N=955).
-GPT-4’s estimation rate of symptom absence (score=0) is very similar to that of self-report, but its estimates were rarely 3 (most severe) and was most commonly 1 for majority of the symptoms.
-Item % Scoring 0 % Scoring 1 % Scoring 2 % Scoring 3
-Self Report GPT-4 Self Report GPT-4 Self Report GPT-4 Self Report GPT-4
-Anhedonia 24.3 28.5 32.5 43.4 24.4 26.5 18.8 1.7
-Depressed Mood 23.8 25.0 32.1 24.6 22.0 44.0 22.2 6.5
-Sleep 25.1 35.2 23.8 54.4 21.1 10.1 30.0 0.4
-Fatigue 15.4 29.3 27.2 50.9 22.7 19.4 34.6 0.4
-Appetite 30.6 57.2 25.9 40.0 19.1 2.7 24.4 0.1
-Worthlessness / Guilt 29.4 43.3 25.4 37.5 23.0 18.3 22.2 0.9
-Difficulty concentrating 28.0 36.0 30.9 58.3 24.1 5.7 16.9 0.0
-Psychomotor Agi./ Ret. 55.5 53.3 23.2 45.3 15.0 1.4 6.4 0.0
-Suicidal ideation 60.3 87.9 21.4 8.8 10.0 2.3 8.3 1.1
+Machine Behavior Analysis of GPT-5
+The decrease in correlations of GPT-5 assessments in comparison to GPT-4 (see Table 1) was explained by its Schema of depression described below.
+Figure 8(c) shows consistently weaker intersymptom relationship compared to self-report.
+The item location analysis also shows larger differences in discriminative properties of symptoms in GPT-5 (Figure 8(c)).
+GPT-5 had a slightly higher rate of symptom identification (Table 9) than GPT-4 for the less observed symptoms (Suicidal Ideation, Psychomotor agitation/retardation, Difficulty Concentration and Appetite).
+However, its overall convergence with self-report was under both explicit and implicit identification showed a small decrease.
+Figure 8: Similarities and differences between GPT-5’s and self-report’s schema of depressive symptoms.
+PHQ9 Item Item Location Rank
+GPT-5 Self-Report Item location diff
+Anhedonia 2 [2, 2] 3 [3, 3] -1
+Depressed Mood 1 [1, 1] 2 [1, 2] -1
+Sleep 6 [4, 6] 4 [4, 5] 2
+Fatigue 3 [3, 3] 1 [1, 2] 2
+Appetite 9 [8, 9] 7 [7, 7] 2
+Worthlessness / Guilt 4 [4, 6] 5 [4, 6] -1
+Difficulty concentrating 5 [4, 6] 6 [4, 6] -1
+Psychomotor 8 [8, 9] 8 [8, 8] 0
+Suicidal Ideation 7 [7, 7.5] 9 [9, 9] 0
+Figure 9: Contrast of explicit & implicit symptom mentions and convergence for GPT-5.
+Pearson Correlation ($r$) between GPT-5 and self-reported scores for all symptoms under conditions where symptoms were identified to be explicit and implicit (N=955) by GPT-5.
+The bar chart in % Explicit column represents the rate of explicit mentions according to GPT-5.
+Higher correlations were observed consistently when symptoms were identified to be explicitly mentioned.
+Symptom Explicit $r$ % Explicit vs. % Implicit Implicit $r$
+Anhedonia .72 41.21% .52
+Depressed Mood .74 88.81% .51
+Sleep .53 12.45% .35
+Fatigue .57 17.36% .40
+Appetite .70 5.65% .20
+Worthlessness / Guilt .56 19.56% .28
+Concentration .46 11.82% .31
+Psychomotor .04 4.60% .11
+Suicidal Ideation .70 12.97% .05
+Mean .56 23.82% .30